### v1

# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities, but still struggle with issues like hallucinations and outdated information. Retrieval-augmented generation (RAG) addresses these issues by grounding LLM outputs in external knowledge with an Information Retrieval (IR) system. Building on this foundation, graph-based RAG systems go a step further by retrieving subgraphs, which preserve the relationships between knowledge entities and provide more comprehensive context. However, graph RAG faces two challenges: (1) Retrieving relevant information introduces irrelevant nodes (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and leads to overly lengthy inputs that hinder efficiency; (2) The representation gap between graph and language during generation with LLMs limits the ability to fully leverage graph structures for enhanced understanding. To address these limitations, we propose Align-GRAG, a novel reasoning-guided dual alignment framework in post-retrieval phrase. It first formulates a subgraph by retrieving nodes and edges. Then an Aligner is proposed to jointly optimizes a graph encoder with LLM-summarized reasoning. It achieves dual alignment of graph node and representation by leveraging KL divergence loss and contrastive loss, facilitating efficient pruning of irrelevant knowledge and establishing a unified semantic space. The Generator integrates the aligned graph data with LLM to produce coherent and accurate answers. Experiments on GraphQA benchmark across three tasks (including common sense reasoning, scene graph understanding, and knowledge graph reasoning) validate the effectiveness of our method. The code will be available upon accepted.
1 Introduction
Recent advancements in natural language processing have demonstrated the remarkable capabilities of large language models (LLMs), such as GPT [1] and Llama [2], in understanding, reasoning, and handling complex tasks [3, 4]. The massive parameterized knowledge encoded within LLMs has enabled them to excel in information retrieval [5, 6, 7]. Despite these advancements, concerns about the interpretability of LLM remain, leading to challenges like hallucinations (fabrication of false or misleading information) [8, 9] and reliance on outdated data [10]. These issues are especially concerning in high-stakes fields such as healthcare [11] and law [12].
Retrieval-augmented generation (RAG) systems have been developed to solve these problems [10]. By integrating external information retrieval (IR) systems, RAG retrieves knowledge from external databases, ensuring that LLMs are equipped with up-to-date and relevant information tailored to the user’s query. By grounding the LLM’s outputs in verifiable, external knowledge, RAG systems significantly improve the reliability and accuracy of content generated by LLMs [13]. Nevertheless, in real-world scenarios, RAG systems often divide long context into independent chunks, overlooking the deeper connections between fragments and lacking a global perspective [14, 15]. Moreover, a significant portion of data inherently exhibits a graph-like structure, such as in recommendation systems [16], the Web, and knowledge graphs [17]. To address these, researchers have proposed graph-based RAG (GRAG) [14, 18, 19, 20] as an enhancement to traditional RAG methods. Instead of retrieving isolated text chunks where information is treated independently of its structural relationships, GRAG retrieves subgraphs from databases [18]. These subgraphs maintain the interconnected relationships between knowledge entities, providing richer context and enabling more accurate, contextually relevant responses.
However, despite its potential, integrating RAG with graphs is non-trivial, presenting two key challenges: (1) Irrelevant Knowledge: The retrieval stage often introduces irrelevant information, hindering the LLM’s attention on useful knowledge (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and the generated inputs become overly lengthy, further causing computational inefficiency. Some traditional RAG approaches [21, 22, 23] aim to tackle this challenge by incorporating a reranker in the post-retrieval stage [13, 10]. For instance, BGE [21] leverages cross-encoder-based architectures to effectively capture interactions between queries and documents, enabling more precise relevance estimation. However, in the GRAG setting, where chunks are typically represented as nodes or triplets, this approach can disrupt structural information and overlook the connections between knowledge. Meanwhile, existing GRAG methods [15] focus primarily on retrieval-stage optimizations (e.g., GraphRAG [14] builds community graphs for long documents to facilitate retrieval and summarization), resulting in relatively shallow similarity that lack deeper semantic matching capabilities. (2) Representation Gap: In the generation stage, when integrating LLMs with graph structure embeddings, such as those generated by Graph Neural Network (GNN), challenge arises due to the representation gap: graph embeddings are designed to capture structural properties, which misaligns with the sequence-based representations used by LLMs. This misalignment limits the ability to fully leverage graph for enhanced reasoning [24, 25]. In the field of LLMs, most existing GRAG methods [20, 26, 27, 28, 29, 30] primarily treat graph data as plain text inputs, lacking a dedicated encoder for structural information, which limits their understanding capabilities. Some methods [19, 31, 18] incorporate graph encoders, but typically rely on straightforward approaches like concatenating graph embeddings with LLM inputs. For instance, G-Retriever [18] uses a projector to map graph embeddings, while GNP [19] applies cross-modality pooling to fuse text and graph representations. However, neither approach explicitly aligns these distinct representation spaces, leaving the underlying gap unresolved. These research gaps present a question: Can we better prune irrelevant nodes while keeping structural information, and effectively encode graph representations for utilization by LLMs?
In this work, we propose Align-GRAG, a novel reasoning-guided dual alignment framework specifically designed for GRAG. During the retrieval stage, we calculate the similarity between nodes and edges to extract the initial subgraph. To address the two key challenges in GRAG, we introduce an innovative graph aligner module for the post-retrieval pharse. We leverage LLM-based summarization ability with well-crafted prompts to generate reasoning chains that bridge information gaps. The aligner achieves dual alignment with reasoning chains as optimization target: (1) Graph Node Alignment, which ranks critical reasoning nodes, pruning irrelevant knowledge; and (2) Graph Representation Alignment, ensuring unified semantic space between graph and language representations via contrastive learning. This dual alignment refines subgraphs to focus on relevant knowledge, enabling the generator to produce accurate, context-aware responses grounded in the graph’s structure. Extensive experiments are conducted on GraphQA benchmark [18], covering tasks such as commonsense reasoning, scene graph understanding, and knowledge graph reasoning, which demonstrate the superior performance of Align-GRAG compared to other baselines and highlight the effectiveness of Aligner module. Overall, this work makes three key contributions:
  • To the best of our knowledge, this is the first work to introduce a reasoning-guided graph alignment framework in the post-retrieval phrase for GRAG.
  • We propose a novel graph aligner module for dual alignment between graph node and representation, allowing effective pruning of irrelevant nodes while learning a unified semantic space.
  • The effectiveness of Align-GRAG was comprehensively validated on three datasets across different tasks, demonstrating superior performance compared to other strong baselines.
2 Preliminaries
This section introduces key concepts and notations, including textual graphs and task formulation.
Textual Graph. A textual graph is a graph where nodes and edges are enriched with textual information, capturing both structural and semantic details. Formally, a textual graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E},\{t_{n}\}_{n\in\mathcal{V}},\{t_{e}\}_{e\in\mathcal{E}})$, where $\mathcal{V}$ and $\mathcal{E}$ are the sets of nodes and edges. $t_{n}\in\mathcal{D}^{L_{n}}$ represents the text associated with a node $n\in\mathcal{V}$, where $\mathcal{D}$ is the vocabulary, and $L_{n}$ is the text length. Similarly, $t_{e}\in\mathcal{D}^{L_{e}}$ is the text for an edge $e\in\mathcal{E}$, with $L_{e}$ as its length.
Task Formulation. This work addresses RAG with textual graph. The goal is to retrieve relevant information from textual graphs and generate accurate responses. Given a query $t_{q}$, a sequence of tokens from $\mathcal{D}$, the model retrieves a subgraph $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ from $\mathcal{G}$ based on the semantic similarity of $t_{n}$, $t_{e}$ with $t_{q}$, where $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$. The retrieved nodes, edges, and texts are refined to improve the input quality for the LLM. During generation, the retrieved subgraph, query $t_{q}$, and prompt $\mathcal{P}$ are provided as input to the LLM, which produces the final output $Y$, grounded in the retrieved knowledge and graph context.
3 Methodology
In this section, we introduce Align-GRAG framework, which is composed of three components: Graph Retriever (Sec. 3.1), Graph Aligner (Sec. 3.2), and Graph Generator (Sec. 3.3) modules. The overall framework pipeline is depicted in Figure 1. In Figure 1, ➀ denotes Graph Retriever, we leverage the similarity between entities, relations, and the query to extract the initial subgraph. In the aligner module (Figure 1-➁), we use a GNN backbone to encode subgraph information. To tackle irrelevant knowledge noise and representation gaps, we introduce an LLM-based summarization technique with tailored prompts. By taking the query-answer pair and subgraph as input, the LLM generates a reasoning chain that links the query to the answer, identifying key intermediate nodes and bridging information gaps. Next, we propose the graph aligner to achieve dual alignment with this summarized reasoning chain: (1) Graph Node Alignment: A KL divergence loss [32] highlights critical nodes and reasoning edges, aligning node-level distributions and pruning irrelevant nodes. (2) Graph Representation Alignment: A contrastive loss with in-batch negative sampling aligns graph and language representations in a shared semantic space. Jointly optimizing these alignments enables effective knowledge pruning and unified representation learning. In the generation stage (Figure 1-➁), the aligned graph embedding is used as a graph token, concatenated with the query and pruned subgraph, and then fed into an LLM for answer generation. By leveraging the aligned graph, generator produces responses that reflect a deep understanding of graph’s structure and its relevance to query.
Figure 1: The Align-GRAG framework includes three components: ➀ Graph Retriever, ➁ Graph Aligner, and ➂ Graph Generator. It first retrieve subgraphs, then trains the aligner with LLM-based summarized reasoning chain. The trained aligner prunes irrelevant nodes, generates graph embeddings, and feeds them into the LLM to produce the final answer.
3.1 Graph Retriever
Existing RAG methodologies are mainly designed for plain text documents or triplets, where information retrieval occurs independently of the graph structure [13, 5, 17]. In retrieval stage, we first utilize an encoder-only language model (e.g., SentenceBERT [33]) to encode textual information, including the query $t_{q}$, text of each node $t_{n}$ and edge $t_{e}$ in graph respectively:
$\boldsymbol{q}=\text{SBERT}(t_{q})\in\mathbb{R}^{d},\boldsymbol{n}=SBERT(t_{n})\in\mathbb{R}^{d},\boldsymbol{e}=SBERT(t_{e})\in\mathbb{R}^{d}$ (1)
Then we compute cosine similarity $\text{sim}(\cdot)$ between query embeddings $\boldsymbol{q}$ and embeddings of nodes $\boldsymbol{n}$ and edges $\boldsymbol{e}$. The top-$k$ nodes and edges are selected as the most relevant entities and relations.
$\mathcal{V}_{k}=\text{argtop}_{k}\big{(}\text{sim}(\boldsymbol{q},\boldsymbol{n})\big{)},\quad\mathcal{E}_{k}=\text{argtop}_{k}\big{(}\text{sim}(\boldsymbol{q},\boldsymbol{e})\big{)}$ (2)
This forms retrieve top-$k$ entities and relation sets, denoted as $\mathcal{G}_{\text{retriever}}$. Inspired by G-retriever [18], we further leverage the Prize-Collecting Steiner Tree algorithm [34] to maintain a controlled graph size, as detailed in Appendix A.
3.2 Graph Aligner
In this section, we introduce the Graph Aligner module. The retrieved subgraph $\mathcal{G}_{\text{Retriever}}$ faces limitations when integrated with LLMs, including irrelevant information from the retriever and misalignment between graph and language representations during generation with LLM. Our Graph Aligner tackles these issues through two key objectives: (1) Aligning node-level distributions to prune irrelevant structures, refining subgraphs based on LLM knowledge preferences; (2) Bridging the gap between graph structures and language descriptions by unifying them in a shared latent space for LLM integration. The illustration is shown in Figure 1-➁. We provide a detailed explanation below.
3.2.1 Summarized Reasoning Chain
To achieve the Aligner module’s optimization goals, our motivation is to leverage crucial reasoning chain related to query-answer pair. This is especially important for multi-hop problems, which require intermediate reasoning steps to bridge information gaps. Incorporating such reasoning significantly improves answer quality, as demonstrated by methods like Chain-of-Thoughts [35] and O1 [36]. To explicitly identify the essential nodes within graph database, we propose an LLM-based summarization technique. By designing well-crafted prompts, we harness the capabilities of strong LLMs to generate comprehensive reasoning chains that logically connect the question to the answer. For example of Table 8 of Appendix G, given a query, “What is the name of the first Harry Potter novel?”, with the LLM-generated prior, we can identify the intermediate and critical node, J.K. Rowling. We believe this can serve as an excellent label. The deeper insights provided through reasoning are subsequently utilized in the dual node and representation alignment process to train the aligner module. More details for summarization prompt are provided in Appendix G.
3.2.2 Node Alignment
The Aligner module utilizes the summarized reasoning chain to identify and align relevant nodes within the graph, effectively pruning redundant nodes and edges by filtering out unrelated information at the node level. Specifically, we employ a GNN, which can be GraphTransformer [37] or GAT [38], to encode the structural information of the graph. The GNN produces node-level embeddings $\boldsymbol{n}_{g}$ based on the input graph:
$\boldsymbol{n}_{g}=\text{GNN}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d}$ (3)
where, $|\mathcal{V}|$ is the number of node and $d$ is feature dim. For the reasoning text, we employ SBERT to encode the textual description into a representation $\boldsymbol{r}_{s}$ that captures its semantic meaning:
$\boldsymbol{r}_{s}=\text{SBERT}(t_{\text{reasoning}})\in\mathbb{R}^{d_{s}}$ (4)
We concatenate the embedding of each node in the subgraph with the query embedding. The concatenated embeddings are then passed through an MLP module, which generates a predicted importance score for each node. Finally, these scores are transformed into probability distributions using the softmax function, producing both predict and reasoning importance scores.
$\boldsymbol{p}_{\text{predict}}=\text{Softmax}(\text{MLP}([\boldsymbol{n}_{g},\boldsymbol{q}^{\text{expanded}}]))\in\mathbb{R}^{|\mathcal{V}|},\boldsymbol{p}_{\text{reasoning}}=\text{Softmax}(\text{sim}(\boldsymbol{r}_{s},\boldsymbol{n}))\in\mathbb{R}^{|\mathcal{V}|}$ (5)
where [,][,][ , ] means concat operation, $\boldsymbol{q}^{\text{expanded}}$ is the query embedding broadcasted across all nodes. To align the node distribution, we minimize the Kullback–Leibler (KL) divergence [32] between predicted probabilities $\boldsymbol{p}_{\text{predict}}$ and the probabilities $\boldsymbol{p}_{\text{reasoning}}$. The KL divergence loss for a subgraph is given by:
$\mathcal{L}_{\text{NA}}=\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\boldsymbol{p}_{\text{reasoning}}(i)\log\frac{\boldsymbol{p}_{\text{reasoning}}(i)}{\boldsymbol{p}_{\text{predict}}(i)}$ (6)
Optimizing $\mathcal{L}_{\text{NA}}$ enables effective alignment of relevant knowledge.
3.2.3 Representation Alignment
To bridge the representation gap between graph structures and language-based descriptions, by treating the text representation derived from reasoning chain as the target label, Aligner module aligns the graph and text embeddings to encourage semantic consistency. We apply a mean pooling operation across the node embeddings $\boldsymbol{n}_{g}$ to obtain a unified graph-level representation $\boldsymbol{r}_{g}$:
$\boldsymbol{r}_{g}=\text{POOL}(\boldsymbol{n}_{g})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\boldsymbol{n}_{g}(v)\in\mathbb{R}^{d}$ (7)
To unify the graph and text representations in a shared semantic space, we apply a contrastive loss with in-batch negative sampling. This loss encourages positive pairs (i.e., graph and text embeddings) to have higher similarity, while pushing apart non-matching pairs. Then, a shared-weight MLP layer is used to map $\boldsymbol{r}_{g}$ and $\boldsymbol{r}_{s}$ to dimension $d_{t}$ of LLM token embeddings:
$\boldsymbol{\hat{r}}_{s}=\text{MLP}(\boldsymbol{r}_{s})\in\mathbb{R}^{d_{t}},\boldsymbol{\hat{r}}_{g}=\text{MLP}(\boldsymbol{r}_{g})\in\mathbb{R}^{d_{t}}$ (8)
The contrastive loss for representation alignment from $\boldsymbol{\hat{r}}_{g}$ to $\boldsymbol{\hat{r}}_{s}$ is defined as:
$\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s})=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{j})/\tau\right)}\right]$ (9)
where, $N$ is the batch size, $(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})$ is the $i$-th positive (graph-text) pair in the batch, $\tau$ is a temperature parameter to control the sharpness. Similarly, we can obtains the loss $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})$ from $\boldsymbol{\hat{r}}_{s}$ to $\boldsymbol{\hat{r}}_{g}$. The final representation alignment loss are obtained by:
$\mathcal{L}_{RA}=1/2(\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})+\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s}))$ (10)
To achieve an optimized Graph Aligner, we perform joint optimization of node and representation Alignment. The total loss for the Graph Aligner is defined as: $\mathcal{L}_{\text{Aligner}}=\mathcal{L}_{\text{RA}}+\mathcal{L}_{\text{NA}}$ The parameters of GNN encoder are jointly optimized using the loss $\mathcal{L}_{\text{Aligner}}$ with a specified training step, which reflects degree of alignment. We evaluate impact of the align degree in Figure 2 and the effectiveness in Section 4.4.
3.2.4 Graph Pruning and Representation Generation
We use the trained Graph Aligner to prune irrelevant nodes, and produce graph representations that are aligned with language. The Aligner performs graph pruning across the entire dataset. To start, we introduce a hyperparameter: the number of seed nodes, denoted as $n_{\text{seed}}$. For the graph’s node-level output, $\boldsymbol{n}_{\text{predict}}$, the top $n_{\text{seed}}$ nodes are selected as seed nodes. Using the first-order neighbors of these seed nodes, we expand the adjacent nodes and edges. This process yields the newly constructed subgraph, denoted as $\mathcal{G}_{\text{Aligner}}=(\mathcal{V}_{\text{Aligner}},\mathcal{E}_{\text{Aligner}})$. What is more, we also utilize the trained Aligner across the entire dataset to produce the graph representation $\boldsymbol{r}_{g}$. The aligned subgraph $\mathcal{G}_{\text{Aligner}}$ serves as the input to LLM, enabling more efficient and accurate generation results. The efficiency of pruning are evaluated in Appendix C.
3.3 Graph Generator
The Aligner module enables efficient representation alignment and pruning of irrelevant knowledge, greatly facilitating integration with LLMs. To generate the final answer, we concatenate the text of graph $\mathcal{G}_{\text{Aligner}}$ with query tokens and input them into the LLM’s TextEmbedder ((The token embedding of LLM itself). This yields text tokens embedding: $\boldsymbol{r}_{t}=\text{TextEmbedder}([t_{q},\mathcal{G}_{\text{Aligner}}])\in\mathbb{R}^{(T_{q}+T_{g})\times d_{t}}$, where $T_{q}$ and $T_{g}$ mean the token length of query and graph. $\boldsymbol{r}_{t}$ is then concatenated with the previously learned graph embedding $\boldsymbol{r}_{g}$. The conditional probability of generating answer $Y$ given the aligned graph $\mathcal{G}_{\text{Aligner}}$ and the query $t_{q}$ is defined as:
$p_{\theta}(Y|t_{q},\mathcal{G}_{\text{Aligner}})=\prod_{i=1}^{m}p_{\theta}(y_{i}|[\text{MLP}(\boldsymbol{r}_{g}),\boldsymbol{r}_{t}],y_{<i})$ (11)
To enhance efficiency of training and deployment, we fine-tune generator using parameter-efficient tuning method, like LoRA [39].
4 Experiments
We conducted experiments on the GraphQA benchmark [18], which includes ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). To ensure a fair comparison, we utilized the same retrieval results obtained via the PCST algorithm across all baselines. More details about datasets, baselines and implementation details are provided in Appendix B.
4.1 Main Results
Table 1: Performance comparison using Llama2-7b [2] and GraphTransformer [37] as backbones and same retrieval settings for all methods. The table reports the mean and standard deviation (std) results across three random seeds. For methods marked with $\dagger$’, we reproduce the results on WebQSP to report both F1 and Accuracy. Results for methods marked with $\ddagger$’ are taken directly from [18] and were not reproduced due to their poor performance. The best results are highlighted in bold, and the second-best results are underlined. ‘Improvement’ represents the gain of Align-GRAG over the second-best baseline. ‘*’ indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. $\uparrow$: higher is better.
Method ExplaGraphs SceneGraphs WebQSP
[…]
Inference-$\ddagger$
Zero-shot 0.5650 0.3974 - 0.4106 -
Zero-CoT [40] 0.5704 0.5260 - 0.5130 -
CoT-BAG [41] 0.5794 0.5680 - 0.3960 -
KAPING [42] 0.6227 0.4375 - 0.5264 -
Raw Fine-tuning
Prompt tuning [43] 0.5763$\pm$0.0243 0.6341$\pm$0.0024 0.2652$\pm$0.0049 0.4807$\pm$0.0055 0.2827$\pm$0.0073
LoRA [39] 0.8538$\pm$0.0353 0.7862$\pm$0.0031 0.4445$\pm$0.0058 0.6505$\pm$0.0068 0.4479$\pm$0.0091
Reranker-based
gte-base [22] 0.8557$\pm$0.0144 0.8556$\pm$0.0095 0.5378$\pm$0.0044 0.7373$\pm$0.0064 0.5251$\pm$0.0052
gte-large [22] 0.8776$\pm$0.0095 0.8592$\pm$0.0074 0.5392$\pm$0.0013 0.7340$\pm$0.0044 0.5374$\pm$0.0038
bge-reranker-base [21] 0.8534$\pm$0.0159 0.8577$\pm$0.0029 0.5323$\pm$0.0052 0.7397$\pm$0.0012 0.5254$\pm$0.0010
bge-reranker-large [21] 0.8612$\pm$0.0184 0.8644$\pm$0.0060 0.5366$\pm$0.0045 0.7391$\pm$0.0093 0.5401$\pm$0.0077
G-RAG [44] 0.8484$\pm$0.0174 0.8474$\pm$0.0147 0.5181$\pm$0.0023 0.7114$\pm$0.0113 0.5080$\pm$0.0041
G-RAG-RL [44] 0.8478$\pm$0.0112 0.8509$\pm$0.0142 0.5291$\pm$0.0066 0.7167$\pm$0.0039 0.5185$\pm$0.0026
GNN-based
[…]
GNP [19] 0.8704$\pm$0.0034 0.8616$\pm$0.0096 0.5369$\pm$0.0049 0.7391$\pm$0.0100 0.5441$\pm$0.0046
[…]
$\text{GRAG}_{PCST}^{\dagger}$ [45] 0.8805$\pm$0.0050 0.8561$\pm$0.0052 0.5355$\pm$0.0049 0.7485$\pm$0.0104 0.5503$\pm$0.0035
Align-GRAG (Ours) 0.8992$\pm$0.0124 0.8804$\pm$0.0106 0.5445$\pm$0.0041 0.7626$\pm$0.0063 0.5700$\pm$0.0039
In this section, we conduct extensive experiments on the ExplaGraphs, SceneGraphs, and WebQSP datasets, comparing the performance of our method with 16 baseline methods. All baselines are implemented using the same retrieved text, and the same backbones, Llama2-7b [2] and GraphTransformer [37]. As shown in Table 1, the overall results demonstrate the effectiveness of Align-GRAG, which achieves performance improvements across all metrics on three datasets. Notably, on WebQSP dataset, Align-GRAG achieves a remarkable 4.76% improvement in Accuracy compared to the second-best method. This improvement can be attributed to the innovative design of Align-GRAG, particularly its alignment-based pruning and optimization strategies for graph structures, showcasing its potential in complex graph reasoning and knowledge-based question answering tasks.
Comparing with Inference-only methods. They rely solely on the reasoning capabilities of LLM without task-specific optimization and, as a result, perform poorly. Specifically, the Zero-shot approach, which does not depend on any carefully constructed prompts, suffers from limited performance due to its inability to leverage task-specific knowledge.
Comparing with raw fine-tuning methods. LoRA significantly outperforms Prompt Tuning and also surpasses inference-only methods. This highlights the effectiveness of task-specific optimization in graph QA scenarios. However, it still lags behind reranker-based and GNN-based methods, indicating that more advanced techniques are required to fully exploit graph structure.
Comparing with reranker-based methods. gte-large achieves the second-best performance on both the ExplaGraphs and WebQSP datasets (F1 metric), outperforming other reranker approaches. However, it still falls significantly short of Align-GRAG. This demonstrates that simply performing node and triple reranking is insufficient to fully capture the graph structural information. Although G-RAG-RL also utilizes a GNN encoder to model graph structures, it rely on similarity with the query as a label, resulting in poor performance. This validates the effectiveness of our approach, which uses reasoning chains summarized by LLM as optimization target. By aligning graph reasoning with LLM-derived reasoning chains, Align-GRAG achieves superior performance, showing that this alignment is crucial for graph-based reasoning tasks.
Comparing with GNN-based methods. G-Retriever ranks second on SceneGraphs, while GNP achieves the second-best Accuracy on WebQSP. However, both methods fall significantly short of Align-GRAG overall. This is because G-Retriever simply concatenates graph embeddings with LLM inputs using a projector, while GNP employs cross-modality pooling to fuse text and graph representations. Despite these efforts, their approaches remain limited. In contrast, Align-GRAG explicitly aligns the representation of graphs with language, effectively bridging the gap between these two spaces and achieving state-of-the-art results.
4.2 Ablation Study
Table 2: Ablation study on different alignment strategy.
Method WebQSP
[…]
Align-GRAG 0.5445 0.7626 0.5700
w/o Representation Alignment 0.5458 0.7586 0.5675
w/o Node Alignment 0.5344 0.7371 0.5339
w/o Both 0.5348 0.7328 0.5216
Random Alignment 0.4617 0.6861 0.4865
We perform an ablation study on WebQSP using three metrics to evaluate how the proposed modules contribute to performance improvement. As illustrated in Table 2, we examine the following variants: (1) w/o Node Alignment: Removes the KL loss and knowledge pruning components. (2) w/o Representation Alignment: Excludes the contrastive loss and the graph embedding input to the LLM. (3) w/o Both: Removes both Node Alignment and Representation Alignment. (4) Random Alignment: Optimizes the aligner module using randomly generated labels. We observe that removing the Node Alignment module significantly lowers all evaluation metrics compared to the full model (Align-GRAG), indicating the effectiveness of node-level alignment optimization and node pruning. In contrast, while excluding the Representation Alignment module also leads to a performance decline, the impact is comparatively less significant. Performance further deteriorates when both modules are removed (w/o Both), highlighting the critical role of dual alignment between node and representation in filtering irrelevant knowledge and bridging the representation gap between the graph and language. The Random Alignment variant performs the worst, as random alignment fails to guide the model effectively. This emphasizes the necessity of meaningful alignment strategies and demonstrates that our LLM-summarized reasoning chain provides a highly useful label.
Table 3: Generalization Analysis on different LLM and GNN backbones.
LLM Backbones Llama-2-7b-hf Llama-2-13b-hf
GNN Method ExplaGraphs WebQSP ExplaGraphs WebQSP
[…]
GT GNP 0.8704 0.7391 0.8880 0.7696
G-Retriever 0.8705 0.7366 0.9115 0.7739
Align-GRAG (Ours) 0.8992 0.7626 0.9241 0.7789
GAT GNP 0.9061 0.7291 0.8989 0.7676
G-Retriever 0.7960 0.7414 0.8953 0.7737
Align-GRAG (Ours) 0.9151 0.7309 0.9151 0.7573
GCN GNP 0.7545 0.7298 0.8682 0.7564
G-Retriever 0.8592 0.7352 0.9007 0.7521
Align-GRAG (Ours) 0.8574 0.7377 0.9152 0.7592
4.3 Generalization Analysis
In this section, we analyze whether the effectiveness of Align-GRAG can generalize across different GNN backbones and LLM sizes. Experiments are conducted on both ExplaGraphs and WebQSP datasets, with various GNN backbones, including GT (GraphTransformer) [37], GAT [38], and GCN [46], as well as LLM backbones of different sizes. As shown in Table 3, our method consistently achieves the best performance across most settings (9 out of 12 experimental settings), demonstrating its strong adaptability. Notably, our method shows the most significant improvement over GNP and G-Retriever when using GT as the backbone, highlighting that GT is particularly well-suited for learning dual alignment of node and representation. On the other hand, we observe that results with the -13B LLM generally outperform those with the -7B LLM, especially when GCN is used as the backbone. This suggests that larger LLMs play a crucial role in effectively learning and reasoning over graph structures.
4.4 Evaluation of Representation Alignment
(a) Align with Query
(b) Align with Summarization
(c) Align with Textualized Graph
Figure 2: Representation Alignment Analysis: The cosine similarity score between graph embeddings and language embeddings (aligned using the aligner module vs. the unaligned setting).
In this section, we evaluate whether the aligner can effectively bridge the representation gap? Specifically, we calculate the Cosine Similarity scores between the graph embeddings and the language embeddings of the query, summarization, and textualized graph, respectively, on test set. The graph embeddings include both the unaligned embeddings and the aligned embeddings (optimized using our contrastive loss). From the results in Figure 2, we observe that as the alignment degree increases, Cosine Similarity scores improve. For instance, similarity with summarization embeddings improves significantly as alignment progresses from 40 to 80 steps. For query and textualized graph embeddings, scores rise sharply between 0-4 steps. This indicates that training effectively reduces the representation gap between graph and language embeddings. After 8 steps, aligned embeddings consistently achieve higher similarity scores then unaligned across all text representations (query, summarization, and textualized graph), demonstrating the effectiveness of our approach. However, as shown in experiment of Appendix E, the alignment degree does not necessarily improve indefinitely. When it reaches a certain level, excessive alignment may over-alter the original graph information, compromising its accuracy.
5 Related Work
Retrieval-Augmented Generation (RAG). RAG [13, 10, 13] has been extensively studied to address the challenges of hallucination and outdated information [8, 9], which often lead to untrustworthy outputs from LLMs [47]. By integrating information retrieval systems, RAG has demonstrated its effectiveness in real-world applications, enhancing reliability of generated responses [48, 49, 50, 51]. The RAG pipeline mainly includes: pre-retrieval (e.g., query rewriting and query expansion), retrieval, post-retrieval (e.g., reranking), and generation [13]. Among these, reranking has emerged as a key technique for refining retrieval results. Initial retrieval [52, 53] often relies on simple similarity scoring. Reranking [22, 21, 44], on the other hand, employs more sophisticated models to reassess and reorder the initially retrieved documents based on their relevance to the query, thereby improving the quality and relevance of retrieved documents. For instance, bge-reranker [21] leverages cross-encoders that perform full attention over the input pair, providing higher accuracy compared to embedding-based models. However, in GRAG scenarios, reranking may disrupt the inherent structural information of graph. Our proposed Align-GRAG effectively performs post-retrieval processing to extract relevant knowledge, while preserves the graph structural information.
Large Language Models on Graph. Graphs, composed of nodes and edges, are essential for modeling real-world relationships across various domains. In recent years, GNN [37, 38, 46, 54] have emerged as a powerful tool for encoding graph structures. With the rise of LLMs and their demonstrated impressive capabilities, there is growing interest in integrating LLMs with graph learning techniques to enhance graph-based tasks [55, 56, 57, 58]. Pioneering works [59, 31, 60] have incorporated GNNs by feeding graph tokens into LLMs. Beyond basic integration, recent studies [61, 62, 63] have focused on tighter architectural fusion, embedding graph neural layers within transformer architectures to enable seamless interaction between graph reasoning and natural language understanding . However, most existing researches [58] focus on tasks like node and graph classification. Our work addresses the challenges of retrieval-based graph QA, leveraging retrieval-augmented methods and graph learning techniques while overcoming the limitations of current frameworks.
Graph RAG. Traditional RAG often struggles to capture structured relational knowledge or global context. Graph Retrieval-Augmented Generation (GRAG) addresses these limitations by leveraging graph-structured knowledge for improved retrieval and reasoning [42, 41, 64, 65]. GRAG utilizes graph databases (e.g., Freebase [66], Wikidata [67]) to retrieve graph elements like triples, paths, or subgraphs [14, 68, 45, 69]. Some approaches [14, 15, 70, 71] focus on constructing large-scale graphs from text. For example, GraphRAG [14] uses LLMs to extract entities, partitions graphs using hierarchical clustering, and generates summaries to support retrieval. Other methods [72, 73, 18, 19] encode graph data with GNNs and integrate it with language models for enhanced reasoning, while some translate natural language queries into logical forms for knowledge graph retrieval [74, 75, 76, 27, 28]. More recently, LLMs have been used as iterative agents for reasoning over knowledge graphs [30, 26, 20, 29]. However, existing GRAG approaches primarily emphasize optimizing retrieval but often lack effective post-retrieval strategies, leading to shallow subgraph-query similarity. To address this, our Align-GRAG introduces a novel dual alignment mechanism that bridges retriever and generator, improving overall performance.
6 Conclusion
This work proposes a novel reasoning-guided dual alignment framework, designed to tackle two key challenges in GRAG: the retrieval of irrelevant nodes causing noisy inputs and the representation gap between graph structures and language models. Align-GRAG introduces an innovative graph aligner module for dual alignment of knowledge and representation, enabling the effective pruning of irrelevant nodes while unifying graph-language representations. Extensive experiments on the GraphQA benchmark demonstrate that Align-GRAG consistently outperforms strong baselines.
Appendix A Prize-Collecting Steiner Tree
To achieve the PCST algorithm, we assign ‘prize’ to nodes and edges based on their similarity to a given query. Relevance is determined through the ranked sets of cosine similarity, $\mathcal{V}_{k}$ and $\mathcal{E}_{k}$, as follows:
$\text{prize}(n)=\begin{cases}k-i,&\text{if }n\in\mathcal{V}_{k}\text{ and }n\text{ is the }i\text{-th ranked node}\\ 0,&\text{otherwise}\end{cases}$ (12)
where $i$ is the rank of $n$ in set $\mathcal{V}_{k}$. Nodes that are not among top $k$ rankings are assigned a prize of zero. The objective of the PCST algorithm is to identify a subgraph that maximizes the total prize of nodes and edges while minimizing the cost:
$\mathcal{G}_{\text{retriever}}=\underset{\begin{subarray}{c}S\subseteq\mathcal{G},\\ S\text{ is connected}\end{subarray}}{\arg\max}\left(\sum_{n\in\mathcal{V}_{S}}\text{prize}(n)+\sum_{e\in\mathcal{E}_{S}}\text{prize}(e)-\text{cost}(S)\right)$ (13)
where $\mathcal{V}_{S}$ and $\mathcal{E}_{S}$ are the sets of nodes and edges in the subgraph $S$, respectively. The cost of constructing the subgraph is defined as $\text{cost}(S)=|\mathcal{E}|\cdot C_{e}$, where $|\mathcal{E}|$ is the number of edges, and $C_{e}$ is a predefined per-edge cost that serves as a regularization parameter to control the subgraph’s size. In this way, we can obtain a preliminary retrieved subgraph $\mathcal{G}_{\text{Retriever}}$.
Appendix B Experimental Settings
Table 4: Statistics of GraphQA Benchmark [18].
Dataset ExplaGraphs SceneGraphs WebQSP
#Training 1,659 59,978 2,826
#Validation 553 19,997 245
#Test 554 20,025 1,628
B.1 Datasets and Metrics.
Following G-retriever [18], we conducted experiments on GraphQA benchmark [18], including ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). The dataset statistics are summarized in Table 4, with a data split of train:validation:test = 6:2:2. The split for WebQSP is derived from the original dataset [77]. For ExplaGraphs and SceneGraphs, we use Accuracy as the evaluation metric, and for WebQSP, we employ F1, Hit@1, and Accuracy to evaluate model performance.
B.2 Baselines.
To evaluate the effectiveness of our proposed method, we compare it with four categories of baseline:
  • Inference-only. This category includes models that leverage frozen LLMs for question answering by using textual graph and query as input. Various prompt designs are employed, including: Zero-shot (answers questions directly based on retrieved information), Zero-CoT [35] (enhances zero-shot reasoning by appending the phrase "Let’s think step by step."), CoT-BAG [41] (adds "Let’s construct a graph with the nodes and edges first." after providing the textual description of graph), and KAPING [42] (retrieves relevant graph triples and prepends them to the input question).
  • Raw Fine-tuning. In this configuration, following [18], we fine-tune the LLM using parameter-efficient tuning methods, without incorporating advanced reranking techniques. It includes popular approaches like Prompt Tuning [43] and LoRA [39].
  • Reranker-based. These baselines use reranking models to refine the ranking of documents before inputting it into the LLM. Including: GTE (General Textual Embedding) [22], with two variants (gte-base, 109M parameters; gte-large, 335M parameters) developed by Alibaba DAMO Academy, trained on large-scale relevance text pairs to enhance retrieval accuracy. BGE (BAAI General Embedding) [21], with two variants (bge-reranker-base, 278M parameters; bge-reranker-large, 560M parameters), are cross-encoders optimized for retrieval-augmented LLMs, offering higher accuracy at the cost of efficiency. G-RAG and G-RAG-RL [44] use GNNs as rerankers, leveraging document connections and semantic information from abstract meaning representation graphs to improve context-aware ranking.
  • GNN-based. This category integrates GNN encoders with LLM embeddings. GraphToken [31] encodes graph structures as explicit prompts for LLMs, boosting graph reasoning performance. GNP [19] introduces a plug-and-play approach using a GNN encoder and cross-modality pooling to enhance LLMs with knowledge graphs. G-Retriever [18] establishes the GraphQA benchmark and employs a RAG method with soft prompting, improving graph-based QA, understanding, and mitigating hallucination. GRAG [45] introduces a divide-and-conquer strategy for efficient textual subgraph retrieval and incorporates both text and graph views into LLMs for graph context-aware generation.
B.3 Implementation Details.
For the retrieval process, we use the same retrieval results obtained through the PCST algorithm across all baselines to ensure a fair comparison. We use GraphTransformer [37], GAT [38], and GCN [46] as GNN encoders, and Llama-2-7b-hf and Llama-2-13b-hf [2] as generators. Both reranker-based and GNN-based methods apply LoRA for fine-tuning, except for GraphToken, whose results are from G-retriever [18]. All methods are compared using the same training hyperparameters where applicable, ensuring fair comparability (e.g., GNN layers and LoRA rank). In the Aligner module, we explore two hyperparameters: alignment degree and the number of seed nodes ($n_{\text{seed}}$), with the analysis shown in Figure 5. To summarize the reasoning chain, we employ Llama-3.1-70B-Instruct [78]. The models were trained on the training set, with optimal hyperparameters and early stopping determined using the validation set. Results are reported on the test set. All experiments were run on two NVIDIA A100 GPUs (80GB each). To implement baselines, for GNP [19], we implemented the Graph Neural Prompting module within our framework, including components such as the GNN encoder, cross-modality pooling, and domain projector. For G-RAG and G-RAG-RL [44], we adopted their ranking approach, combining cross-entropy loss with pairwise ranking loss. However, since treating documents as nodes was infeasible in our case, we instead treated entities as nodes and employed the same GNN encoder as our method. We implemented GTE [22] and BGE-reranker [21] using their official open-source models. Nodes and triples are ranked by query similarity and, as in our method, fed into the LLM for generation. For the $\text{GRAG}_{PCST}$ model [45], we reproduced its experiments. However, to ensure a fair evaluation, we standardized the retrieval process by using the PCST method for retrieval. This allowed us to directly compare it with their graph encoder approach.
Appendix C Evaluating the Efficiency after Node Pruning
Table 5: Evaluation of Efficiency. $\uparrow$ indicates results better than PCST, while $\downarrow$ indicates worse results. (xx%) represents the percentage of tokens relative to PCST.
Method # Tokens Infer Time Hit@1
Non-Retriever 100626.28 OOM -
BM25 [52] 2569.57 17:24 min 0.4394
PCST [34] 2576.66 17:28 min 0.4502
PCST
   w/ $\text{Aligner}_{n_{\text{seed}}=4}$ 496.54 (19.27%) 3:31 min 0.4299$\downarrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=6}$ 698.54 (27.11%) 4:45 min 0.4527$\uparrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=8}$ 905.52 (35.14%) 6:55 min 0.4699$\uparrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=10}$ 1120.75 (43.5%) 8:27 min 0.4785$\uparrow$
   w/ $\text{Aligner}_{n_{\text{seed}}=15}$ 1546.90 (60.04%) 11:50 min 0.4914$\uparrow$
In this section, we evaluate the efficiency improvements brought by node alignment and node pruning in Aligner module. Specifically, we compare different retrieval methods in terms of average token consumption, average inference time, and Hit@1 scores on the test set. The w/ Aligner method is built on the initial retrieval results from PCST, where different seed nodes are selected to perform node pruning. The experiments were conducted on two Nvidia A100 80G GPUs in an inference-only setting. To fully utilize GPU resources and memory, we adjusted the batch size to be as large as possible under different average token conditions. For comparison, the BM25 method retrieves triples and ensures that the token consumption is roughly similar to that of PCST. From Table 5, we observe that when the number of seed nodes ($n_{\text{seed}}$) is set to 6, our Aligner method achieves performance comparable to PCST while utilizing 70% fewer tokens. This demonstrates the method’s ability to effectively prune irrelevant knowledge while preserving the essential information. However, when $n_{\text{seed}}$ is too small (e.g., $n_{\text{seed}}=4$), performance drops significantly, suggesting that some critical information may be pruned during the process. On the other hand, increasing the number of seed nodes to 15 (utilizing 60.04% of the tokens) leads to performance that significantly surpasses PCST. This highlights the strength of our method in efficiently selecting useful knowledge and confirms that, in longer-context scenarios, removing irrelevant information enhances overall results. Additionally, we find that BM25 performs worse than PCST when using a similar number of tokens. This suggests that the PCST retrieval method is better suited for graph databases, as it more effectively captures the connections between knowledge.
Appendix D Impact of Seed Nodes
In this section, we analyze the impact of the number of seed nodes ($n_{\text{seed}}$) on model performance. The experiments are conducted on the WebQSP dataset, where we evaluate the Hit@1 and Accuracy metrics. From the experiments in Figure 3, we observe that the Hit@1 and Accuracy performance peaks when the number of seed nodes is set to 25. Beyond this point (from 25 to 30), the performance starts to decline. This indicates that including too many nodes introduces a significant amount of irrelevant knowledge, which negatively impacts the model. Our pruning strategy effectively eliminates irrelevant knowledge to enhance model performance. On the other hand, when the number of seed nodes is as low as 5, the performance is considerably poor. This suggests that excessive pruning removes crucial knowledge, which is detrimental to performance. This highlights a trade-off: pruning reduces noise and improves performance, but over-pruning leads to the loss of essential knowledge.
Figure 3: Hyperparameters Analysis of the Number of seed nodes.
Appendix E Impact of Align Degree
This section examines how the Align Degree (number of training steps for Aligner module) influences model performance. As shown in Figure 4, we evaluate the Hit@1 and Accuracy metrics on the WebQSP dataset. From the experimental curve of Align Degree, we observe that the Hit@1 and Accuracy metrics peak at around 60 epochs before declining. This indicates that, as training progresses, bridging the representation gap between the graph and language helps the LLM better understand graph data. However, excessive training may lead to overfitting, which disrupts graph information and ultimately causes a drop in performance. This suggests that there is an optimal point in training where the representation alignment is most effective, and training beyond this point can be detrimental.
Figure 4: Hyperparameters Analysis of Align degree.
Appendix F Impact of Top K retrieval
Figure 5 illustrates the impact of varying the Top K retrieval of entities and relations on model performance across Accuracy-score, F1-score, and Hit@1-score. By analyzing the trends in the graphs, we can derive insights into the effect of Top K on model performance. All three metrics show a similar trend: performance improves as K increases, peaks at K = 10, and then declines. This suggests that K = 10 strikes the optimal balance between retrieving relevant results and avoiding noise. Smaller K values may miss relevant information, while larger K values dilute relevance, reducing precision and retrieval quality. These findings highlight the importance of selecting an appropriate K value to maximize performance in retrieval-based systems.
Figure 5: Effect of Top K retrieval.
Appendix G Evaluating the Quality of Summarization Prompts
In this section, we conduct a fine-grained analysis of the quality of summarization for reasoning chain. This analysis is divided into three aspects. First, a case study is used to gain a deeper understanding of the forms of summarization. Second, we evaluate Relevance and Faithfulness to assess whether the generated chain are aligned with the questions and answers. Finally, we compare the QA performance of poorer-quality reasoning chain to highlight their impact.
G.1 Evaluation of Relevance and Faithfulness
To assess the quality of generated reasoning chain in terms of their relevance to question-answer context and their faithfulness to provided contextual information, we adopted LLM-as-a-judge framework [79]. Specifically, we utilized GPT-4o [1] to perform the evaluation. A total of 100 samples were randomly selected for this analysis. The evaluation was conducted using a carefully designed prompt, as illustrated in Figure 6. The results of the experiment are presented in Table 6. The analysis reveals that the summarized reasoning chain exhibit a high degree of relevance and faithfulness. Specifically, the summaries achieved a relevance score of 90%, indicating that the majority of the generated content aligns well with the given question-answer context. A faithfulness score of 87% demonstrates that the summaries adhere largely to the factual information provided by the context, with minimal hallucination.
Relevance and Faithfulness Prompt Evaluation of Relevance: Evaluate the relevance of the Reasoning Chain in answering the QUESTION. A relevant Reasoning Chain contains information that helps answer the question, even if partially. Return one of the following labels: ’Relevant’, or ’Irrelevant’ without any additional response. Evaluation of Faithfulness: Evaluate the following Reasoning Chain for faithfulness in answering the QUESTION. A faithful response should include information that helps answer the question, even if partially, avoid inventing new details, and not contradict the context. Return one of the following labels: ’Faithful’ or ’Not Faithful’ without any additional response.
Figure 6: Evaluation of Relevance and Faithfulness Prompt.
Table 6: Evaluating Relevance and Faithfulness of Reasoning Chain.
Metrics WebQSP
Relevance 90%
Faithfulness 87%
G.2 Quantitative comparison
In this experiment, we evaluate the impact of reasoning quality on QA performance by comparing two approaches: $\text{Reasoning}_{\text{w/ answer}}$, where both the question and answer are included during reasoning chain generation, and $\text{Reasoning}_{\text{w/o answer}}$, where only the question is used for reasoning chain generation. The results in Table 7 demonstrate that incorporating the answer during summary generation ($\text{Reasoning}_{\text{w/ answer}}$) consistently outperforms the approach that excludes the answer ($\text{Reasoning}_{\text{w/o answer}}$) across all evaluation metrics. Specifically, the F1 score of $\text{Reasoning}_{\text{w/ answer}}$ is 0.5445, which is higher than the 0.5209 achieved by $\text{Reasoning}_{\text{w/o answer}}$, indicating that including the answer improves the balance of precision and recall in QA tasks. Similarly, the Hit@1 value increases from 0.7444 to 0.7626, showing that the inclusion of the answer helps the model identify correct nodes more accurately in graph alignment. These findings confirm that using the answer during summary generation leads to higher-quality summaries, which in turn positively impacts QA performance, whereas excluding the answer results in inferior summaries and reduced performance.
Table 7: Quantitative Comparison.
Method WebQSP
[…]
$\text{Reasoning}_{\text{w/ answer}}$ 0.5445 0.7626 0.5700
$\text{Reasoning}_{\text{w/o answer}}$ 0.5209 0.7444 0.5409
G.3 Case study
The goal of this case study is to analyze the effectiveness of summarization in identifying intermediate nodes that are critical for reasoning chains, as shown in 8. By leveraging well-crafted prompts and a graph database, we aim to demonstrate how the summaries enable the logical connection from the query to the final answer. Specifically, we assess the ability of the summarization to extract key intermediate nodes, such as ‘J.K. Rowling’, which play a pivotal role in reasoning. Without summarization, the raw graph database alone is insufficient to provide the answer directly.
Table 8: The case for summarizing the reasoning chain. The text from SUMMARIZE_PROMPT, Question, Answer, and Graph DataBase are concatenated as input to generate the Response from LLM. We can see that LLM’s summarization accurately identified middle and critical node in graph, ‘J.K. Rowling’, for answer.
SUMMARIZE_PROMPT: You are a helpful assistant responsible for generating a comprehensive summary of the data provided below. Given question and answer, and related graph data base. Please concatenate all of these into a single, comprehensive description. The description should logically connect the question to the answer. Make sure to include information collected from all descriptions.
Question: what is the name of the first harry potter novel?
Answer: harry potter and the philosopher’s stone
Graph DataBase:
node_id,node_attr
0,harry potter and the chamber of secrets
1,harry potter and the philosopher’s stone
3,j. k. rowling
4,the complete idiot’s guide to the world of harry potter (complete idiot’s guide to)
7,harry potter and the half-blood prince
9,harry potter and the prisoner of azkaban
11,harry potter and the goblet of fire
16,harry potter
24,harry potter and the deathly hallows
43,harry potter and the order of the phoenix
46,harry potter and the deathly hallows: part i
57,fiction
59,harry potter literary series
76,professor severus snape
91,harry potter fanbase
98,fantasy
127,harry potter and the deathly hallows (book 7)
src,edge_attr,dst
16,freebase.equivalent_topic.equivalent_domain,91
91,freebase.domain_profile.featured_views,806
91,freebase.domain_profile.featured_views,790
91,freebase.domain_profile.featured_views,759
199,book.written_work.subjects,455
59,book.book_subject.works,199
3,book.author.works_written,670
24,media_common.adapted_work.adaptations,46
59,book.book_subject.works,371
59,book.book_subject.works,305
0,book.book.characters,325
7,book.book.characters,178
43,book.book.characters,16
24,book.book_edition.book,24
24,book.book.genre,57
24,book.book_edition.book,24
9,book.book.genre,98
190,book.book_edition.book,24
1,book.book.genre,224
24,book.book_edition.book,24
7,book.book.genre,98
59,book.literary_series.fictional_universe,16
478,book.book_edition.book,24
Summarized Reasoning Chain: J.K. Rowling wrote the first Harry Potter novel, “Harry Potter and the Philosopher’s Stone.”
Appendix H LLM Prompts
We provide the Summarization Prompt and Generator Prompt in Figure 7 and Figure 8. Summarization prompt is designed to generate a comprehensive summary by combining a given question, its answer, and related graph data. The output should logically connect the question to the answer while incorporating relevant information from the textualized graph data. The generator prompt varies by dataset. For WebQSP and SceneGraphs, it requires answering a question based on textualized graph data. For ExplaGraphs, it determines whether two arguments support or counter each other, providing the answer as either ‘support’ or ‘counter’.
Summarization Prompt You are a helpful assistant responsible for generating a comprehensive summary of the data provided below. Given question and answer, and related graph data base. Please concatenate all of these into a single, comprehensive description. The description should logically connect the question to the answer. Make sure to include information collected from all descriptions. Question: {question} Answer: {Answer} Graph DataBase: {Textualized Graph}
Figure 7: Prompt for Summarization.
Generator Prompt Prompt for WebQSP and SceneGraphs datasets: Textualized Graph: {Textualized Graph}. Please answer the given question. Question: {question} Answer: Prompt for ExplaGraphs dataset: Textualized Graph: {Textualized Graph}. Argument 1: {arg1} Argument 2: {arg1} Question: Do argument 1 and argument 2 support or counter each other? Answer in one word in the form of ‘support’ or ‘counter’. Answer:
Figure 8: Prompt for Generator.
Appendix I Limitations
Despite the promising results demonstrated by Align-GRAG, this work has certain limitations. Due to resource constraints, we were unable to conduct experiments on larger LLMs, leaving the effectiveness of the proposed alignment approach on more powerful models uncertain. Additionally, since our method requires the generation and utilization of graph embeddings, it cannot be directly implemented on closed-source models, such as GPT-4, which restricts access to internal embedding representations. These limitations highlight potential areas for future exploration, such as validating the scalability of the approach with state-of-the-art LLMs and developing techniques to adapt Align-GRAG for closed-source environments.