### v2

# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities, but still struggle with issues like hallucinations and outdated information. Retrieval-augmented generation (RAG) addresses these issues by grounding LLM outputs in external knowledge with an Information Retrieval (IR) system. Building on this foundation, graph-based RAG systems go a step further by retrieving subgraphs, which preserve the relationships between knowledge entities and provide more comprehensive context. However, graph RAG faces two challenges: (1) Retrieving relevant information introduces irrelevant nodes (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and leads to overly lengthy inputs that hinder efficiency; (2) The representation gap between graph and language during generation with LLMs limits the ability to fully leverage graph structures for enhanced understanding. To address these limitations, we propose Align-GRAG, a novel reasoning-guided dual alignment framework in post-retrieval phrase. It first formulates a subgraph by retrieving nodes and edges. Then an Aligner is proposed to jointly optimize a graph encoder with an LLM-summarized reasoning chain. It achieves dual alignment of graph node and representation by leveraging KL divergence loss and contrastive loss, facilitating efficient pruning of irrelevant knowledge and establishing a unified semantic space. The Generator integrates the aligned graph data with LLM to produce coherent and accurate answers. Experiments on the GraphQA benchmark across three tasks (including common sense reasoning, scene graph understanding, and knowledge graph reasoning) validate the effectiveness of our method. The codes are available in this repository [^1] https://anonymous.4open.science/r/Align-GRAG-F3D8/.
Retrieval-Augmented Generation, Large Language Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
Recent advancements in natural language processing have demonstrated the remarkable capabilities of large language models (LLMs), such as GPT (OpenAI, 2024) and Llama (Touvron et al., 2023), in understanding, reasoning, and handling complex tasks (Zhao et al., 2023; Xu et al., 2024a). The massive parameterized knowledge encoded within LLMs has enabled them to excel in information retrieval (Zhu et al., 2023; AlKhamissi et al., 2022; Petroni et al., 2019). Despite these advancements, concerns about the interpretability of LLM remain, leading to challenges like hallucinations (fabrication of false or misleading information) (Huang et al., 2023; Xu et al., 2024c) and reliance on outdated data (Gao et al., 2023). These issues are especially concerning in high-stakes fields such as healthcare (He et al., 2023) and law (Lai et al., 2024).
Retrieval-augmented generation (RAG) systems have been developed to solve these problems (Gao et al., 2023). By integrating external information retrieval (IR) systems, RAG retrieves knowledge from external databases, ensuring that LLMs are equipped with up-to-date and relevant information tailored to the user’s query. By grounding the LLM’s outputs in verifiable, external knowledge, RAG systems significantly improve the reliability and accuracy of content generated by LLMs (Fan et al., 2024). Nevertheless, in real-world scenarios, RAG systems often divide long context into independent chunks, overlooking the deeper connections between fragments and lacking a global perspective (Edge et al., 2024; Guo et al., 2024). Moreover, a significant portion of data inherently exhibits a graph-like structure, such as in recommendation systems (Guo et al., 2020), the Web, and knowledge graphs (Pan et al., 2024). To address these, researchers have proposed graph-based RAG (GRAG) (Edge et al., 2024; He et al., 2024; Tian et al., 2024; Sun et al., [n. d.]) as an enhancement to traditional RAG methods. Instead of retrieving isolated text chunks where information is treated independently of its structural relationships, GRAG retrieves subgraphs from databases (He et al., 2024). These subgraphs maintain the interconnected relationships between knowledge entities, providing richer context and enabling more accurate, contextually relevant responses.
Figure 1. Comparison of Align-GRAG with traditional RAG (including post-retrieval methods such as reranker). Left: query and textual graph (entities and relations). Middle (a): Traditional RAG retrieves top‑k passages and may use a reranker, but still suffers from noisy/irrelevant evidence and loses the underlying graph structure, which hurts answer grounding. Right (b): Align‑GRAG introduces a graph aligner that jointly performs node alignment and representation alignment, preserving structural relations and filtering irrelevant content, showcasing its effectiveness in graph RAG scenarios.
However, despite its potential, integrating RAG with graphs is non-trivial, presenting two key challenges: (1) Irrelevant Knowledge: The retrieval stage often introduces irrelevant information, hindering the LLM’s attention on useful knowledge (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and the generated inputs become overly lengthy, further causing computational inefficiency. As shown in Figure 1(a), some traditional RAG approaches (Xiao et al., 2024; Li et al., 2023b; Jia et al., 2024) aim to tackle this challenge by incorporating a reranker in the post-retrieval stage (Fan et al., 2024; Gao et al., 2023). For instance, BGE (Xiao et al., 2024) leverages cross-encoder-based architectures to effectively capture interactions between queries and documents, enabling more precise relevance estimation. However, in the GRAG setting, where chunks are typically represented as nodes or triplets, this approach can disrupt structural information and overlook the connections between knowledge. Meanwhile, existing GRAG methods (Guo et al., 2024) focus primarily on retrieval-stage optimizations (e.g., GraphRAG (Edge et al., 2024) builds community graphs for long documents to facilitate retrieval and summarization), resulting in relatively shallow similarity that lacks deeper semantic matching capabilities. (2) Representation Gap: In the generation stage, when integrating LLMs with graph structure embeddings, such as those generated by Graph Neural Network (GNN), a challenge arises due to the representation gap: graph embeddings are designed to capture structural properties, which misalign with the sequence-based representations used by LLMs. This misalignment limits the ability to fully leverage the graph for enhanced reasoning (Liu et al., 2023; Zhao et al., 2022). In the field of LLMs, most existing GRAG methods (Sun et al., [n. d.]; LUO et al., [n. d.]; Xu et al., 2025; Hu et al., 2022; Chen et al., 2024; Jiang et al., 2024) primarily treat graph data as plain text inputs, lacking a dedicated encoder for structural information, which limits their understanding capabilities. Some methods (Tian et al., 2024; Perozzi et al., 2024; He et al., 2024) incorporate graph encoders, but typically rely on straightforward approaches like concatenating graph embeddings with LLM inputs. For instance, G-Retriever (He et al., 2024) uses a projector to map graph embeddings, while GNP (Tian et al., 2024) applies cross-modality pooling to fuse text and graph representations. However, neither approach explicitly aligns these distinct representation spaces, leaving the underlying gap unresolved. These research gaps present a question: Can we better prune irrelevant nodes while keeping structural information, and effectively encode graph representations for utilization by LLMs?
In this work, we propose Align-GRAG, a novel reasoning-guided dual alignment framework specifically designed for GRAG. During the retrieval stage, we calculate the similarity between nodes and edges to extract the initial subgraph. To address the two key challenges in GRAG, we introduce an innovative graph aligner module for the post-retrieval phase, as shown in Figure 1(b). We leverage LLM-based summarization ability with well-crafted prompts to generate reasoning chains that bridge information gaps. The aligner achieves dual alignment with reasoning chains as optimization target: (1) Graph Node Alignment, which ranks critical reasoning nodes, pruning irrelevant knowledge; and (2) Graph Representation Alignment, ensuring unified semantic space between graph and language representations via contrastive learning. This dual alignment refines subgraphs to focus on relevant knowledge, enabling the generator to produce accurate, context-aware responses grounded in the graph’s structure. Extensive experiments are conducted on GraphQA benchmark (He et al., 2024), covering tasks such as commonsense reasoning, scene graph understanding, and knowledge graph reasoning, which demonstrate the superior performance of Align-GRAG compared to other baselines and highlight the effectiveness of Aligner module. Overall, this work makes three key contributions:
  • To the best of our knowledge, this is the first work to introduce a reasoning-guided graph alignment framework in the post-retrieval phrase for GRAG.
  • We propose a novel graph aligner module for dual alignment between graph node and representation, allowing effective pruning of irrelevant nodes while learning a unified semantic space.
  • The effectiveness of Align-GRAG was comprehensively validated on three datasets across different tasks, demonstrating superior performance compared to other strong baselines.
2. Related Work
2.1. Retrieval-Augmented Generation (RAG)
RAG (Fan et al., 2024; Gao et al., 2023; Fan et al., 2024) has been extensively studied to address the challenges of hallucination and outdated information (Huang et al., 2023; Xu et al., 2024c), which often lead to untrustworthy outputs from LLMs (Lewis et al., 2020). By integrating information retrieval systems, RAG has demonstrated its effectiveness in real-world applications, enhancing reliability of generated responses (Salemi and Zamani, 2024; Zamani and Bendersky, 2024; Yang et al., 2024; Sudhi et al., 2024). The RAG pipeline mainly includes: pre-retrieval (e.g., query rewriting and query expansion), retrieval, post-retrieval (e.g., reranking), and generation (Fan et al., 2024). Among these, reranking has emerged as a key technique for refining retrieval results. Initial retrieval (Robertson et al., 2009; Izacard et al., 2021) often relies on simple similarity scoring. Reranking (Li et al., 2023b; Xiao et al., 2024; Dong et al., 2024), on the other hand, employs more sophisticated models to reassess and reorder the initially retrieved documents based on their relevance to the query, thereby improving the quality and relevance of retrieved documents. For instance, bge-reranker (Xiao et al., 2024) leverages cross-encoders that perform full attention over the input pair, providing higher accuracy compared to embedding-based models. However, in GRAG scenarios, reranking may disrupt the inherent structural information of graph. Our proposed Align-GRAG effectively performs post-retrieval processing to extract relevant knowledge, while preserves the graph structural information.
2.2. Large Language Models on Graph
Graphs, composed of nodes and edges, are essential for modeling real-world relationships across various domains. In recent years, GNN (Shi et al., 2020; Veličković et al., 2018; Kipf and Welling, 2016; Wu et al., 2020) have emerged as a powerful tool for encoding graph structures. With the rise of LLMs and their demonstrated impressive capabilities, there is growing interest in integrating LLMs with graph learning techniques to enhance graph-based tasks (Huang et al., 2024b; Jin et al., 2024; Li et al., 2023a; Ren et al., 2024). Pioneering works (Tang et al., 2024; Perozzi et al., 2024; Chai et al., 2023) have incorporated GNNs by feeding graph tokens into LLMs. Beyond basic integration, recent studies (Qin et al., 2023; Zhu et al., 2024; Huang et al., 2024a) have focused on tighter architectural fusion, embedding graph neural layers within transformer architectures to enable seamless interaction between graph reasoning and natural language understanding . However, most existing researches (Ren et al., 2024) focus on tasks like node and graph classification. Our work addresses the challenges of retrieval-based graph QA, leveraging retrieval-augmented methods and graph learning techniques while overcoming the limitations of current frameworks.
2.3. Graph RAG
Traditional RAG often struggles to capture structured relational knowledge or global context. Graph Retrieval-Augmented Generation (GRAG) addresses these limitations by leveraging graph-structured knowledge for improved retrieval and reasoning (Baek et al., 2023; Wang et al., 2024; Han et al., 2024; Xu et al., 2024b; Zhang et al., 2025). GRAG utilizes graph databases (e.g., Freebase (Bollacker et al., 2008), Wikidata (Vrandečić and Krötzsch, 2014)) to retrieve graph elements like triples, paths, or subgraphs (Edge et al., 2024; Mavromatis and Karypis, 2024; Hu et al., 2024; Wu et al., 2024; Xiang et al., 2025; Xiao et al., 2025). Some approaches (Edge et al., 2024; Guo et al., 2024; Fan et al., 2025; Liang et al., 2024) focus on constructing large-scale graphs from text. For example, GraphRAG (Edge et al., 2024) uses LLMs to extract entities, partitions graphs using hierarchical clustering, and generates summaries to support retrieval. Other methods (Yasunaga et al., 2021; Jiang et al., 2022; He et al., 2024; Tian et al., 2024) encode graph data with GNNs and integrate it with language models for enhanced reasoning, while some translate natural language queries into logical forms for knowledge graph retrieval (Zhang et al., 2023; Yu et al., 2022; Luo et al., 2023; Xu et al., 2025; Hu et al., 2022). More recently, LLMs have been used as iterative agents for reasoning over knowledge graphs (Jiang et al., 2024; LUO et al., [n. d.]; Sun et al., [n. d.]; Chen et al., 2024; Zhang et al., 2024). However, existing GRAG approaches primarily emphasize optimizing retrieval but often lack effective post-retrieval strategies, leading to shallow subgraph-query similarity. To address this, our Align-GRAG introduces a novel dual alignment mechanism that bridges retriever and generator, improving overall performance.
3. Preliminaries
This section presents relevant concepts and notations, including the definition of textual graphs and the task formalization.
Textual Graph. A textual graph refers to a graph where the nodes and edges are supplemented with textual information, enabling the representation of both structural relationships and semantic content. Formally, a textual graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\{t_{n}\}_{n\in\mathcal{V}},\{t_{e}\}_{e\in\mathcal{E}})$, where, $\mathcal{V}$ and $\mathcal{E}$ represent the sets of nodes and edges, respectively. $t_{n}\in\mathcal{D}^{L_{n}}$ denotes the sequential text associated with a node $n\in\mathcal{V}$, where $\mathcal{D}$ is the vocabulary set, and $L_{n}$ is the length of the text associated with the node. Similarly, $t_{e}\in\mathcal{D}^{L_{e}}$ represents the sequential text associated with an edge $e\in\mathcal{E}$, and $L_{e}$ denotes the length of the text associated with the edge.
Task Formulation. This work focuses on RAG with textual graph databases. The goal is to retrieve relevant information from the structured nature of textual graphs and generate coherent, contextually accurate responses. Formally, a given query $t_{q}=[x_{1},...,x_{n}]$, represented as a sequence of tokens from the vocabulary $\mathcal{D}$. The model retrieves relevant subgraphs $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ from $\mathcal{G}$ by computing the semantic similarity between the query $t_{q}$ and the node and edge texts $t_{n}$ and $t_{e}$. where $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$. The retrieval process identifies relevant nodes, edges, and their associated texts in the context of the input question $t_{q}$. In the post-retrieval stage, we focus on refining the retrieved information to enhance the quality of input provided to LLM, ensuring more accurate and contextually relevant outputs. In the generation stage, the retrieved subgraph, along with the query $t_{q}$ and prompt $\mathcal{P}$, is provided as input to LLM. The LLM generates the final output $Y=[y_{1},...,y_{m}]$, grounded in the retrieved knowledge and the contextual understanding of the graph structure.
4. Methodology
In this section, we introduce the Align-GRAG framework, which is composed of three components: Graph Retriever (Sec. 4.1), Graph Aligner (Sec. 4.2), and Graph Generator (Sec. 4.3) modules. The overall framework pipeline is depicted in Figure 2. In Figure 1, ➀ denotes Graph Retriever, we leverage the similarity between entities, relations, and the query to extract the initial subgraph. In the aligner module (Figure 2-➁), we use a GNN backbone to encode subgraph information. To tackle irrelevant knowledge noise and representation gaps, we introduce an LLM-based summarization technique with tailored prompts. By taking the query-answer pair and subgraph as input, the LLM generates a reasoning chain that links the query to the answer, identifying key intermediate nodes and bridging information gaps. Next, we propose the graph aligner to achieve dual alignment with this summarized reasoning chain: (1) Graph Node Alignment: A KL divergence loss (Kullback and Leibler, 1951) is utilized to highlight critical nodes and reasoning edges, aligning node-level distributions and pruning irrelevant nodes. (2) Graph Representation Alignment: A contrastive loss with in-batch negative sampling is further designed to align graph and language representations in a shared semantic space. Jointly optimizing these alignments enables effective knowledge pruning and unified representation learning. In the generation stage (Figure 2-➁), the aligned graph embedding is used as a graph token, concatenated with the query and pruned subgraph, and then fed into an LLM for answer generation. By leveraging the aligned graph, generator produces responses that reflect a deep understanding of graph’s structure and its relevance to query.
Figure 2. The Align-GRAG framework includes three components: ➀ Graph Retriever, ➁ Graph Aligner, and ➂ Graph Generator. It first retrieves subgraphs, then trains the aligner with an LLM-based summarized reasoning chain. The trained aligner prunes irrelevant nodes, generates graph embeddings, and feeds them into the LLM to produce the final answer.
4.1. Graph Retriever
Existing RAG methodologies are mainly designed for plain text documents or triplets, where information retrieval occurs independently of the graph structure (Fan et al., 2024; Zhu et al., 2023; Pan et al., 2024). In the retrieval stage, we first utilize an encoder-only language model (e.g., SentenceBERT (Reimers and Gurevych, 2020)) to encode textual information, including the query $t_{q}$, text of each node $t_{n}$ and edge $t_{e}$ in the graph, respectively:
(1) $\boldsymbol{q}$ $=\text{SBERT}(t_{q})\in\mathbb{R}^{d},$
(2) $\boldsymbol{n}$ $=\text{SBERT}(t_{n})\in\mathbb{R}^{d},$
(3) $\boldsymbol{e}$ $=\text{SBERT}(t_{e})\in\mathbb{R}^{d}$
Then we compute cosine similarity $\text{sim}(\cdot)$ between query embeddings $\boldsymbol{q}$ and embeddings of nodes $\boldsymbol{n}$ and edges $\boldsymbol{e}$. The top-$k$ nodes and edges are selected as the most relevant entities and relations.
(4) $\mathcal{V}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{n}\in\mathcal{V}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{n})\big),$
(5) $\mathcal{E}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{e}\in\mathcal{E}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{e})\big)$
This forms top-$k$ entities and relation sets, denoted as $\mathcal{G}_{\text{retriever}}$. Inspired by He et al. (2024), we further leverage the Prize-Collecting Steiner Tree algorithm (PCST) (Bienstock et al., 1993) to maintain a controlled graph size.
To achieve the PCST algorithm, we assign “prize” to nodes and edges based on their similarity to a given query. Relevance is determined through the ranked sets of cosine similarity, $\mathcal{V}_{k}$ and $\mathcal{E}_{k}$, as follows:
[…]
where $i$ is the rank of $n$ in set $\mathcal{V}_{k}$. Nodes that are not among top $k$ rankings are assigned a prize of zero. The objective of the PCST algorithm is to identify a subgraph that maximizes the total prize of nodes and edges while minimizing the cost:
(7) $\mathcal{G}_{\text{retriever}}=\underset{\begin{subarray}{c}S\subseteq\mathcal{G},\\ S\text{ is connected}\end{subarray}}{\arg\max}\left(\sum_{n\in\mathcal{V}_{S}}\text{prize}(n)+\sum_{e\in\mathcal{E}_{S}}\text{prize}(e)-\text{cost}(S)\right)$
where $\mathcal{V}_{S}$ and $\mathcal{E}_{S}$ are the sets of nodes and edges in the subgraph $S$, respectively. The cost of constructing the subgraph is defined as $\text{cost}(S)=|\mathcal{E}|\cdot C_{e}$, where $|\mathcal{E}|$ is the number of edges, and $C_{e}$ is a predefined per-edge cost that serves as a regularization parameter to control the subgraph’s size. In this way, we can obtain a preliminary retrieved subgraph $\mathcal{G}_{\text{Retriever}}$.
4.2. Graph Aligner
In this section, we introduce the Graph Aligner module. The retrieved subgraph $\mathcal{G}_{\text{Retriever}}$ faces limitations when integrated with LLMs, including irrelevant information from the retriever and misalignment between graph and language representations during generation with LLM. Our Graph Aligner tackles these issues through two key objectives: (1) Aligning node-level distributions to prune irrelevant structures, refining subgraphs based on LLM knowledge preferences; (2) Bridging the gap between graph structures and language descriptions by unifying them in a shared latent space for LLM integration. The illustration is shown in Figure 2-➁. We provide a detailed explanation below.
4.2.1. Summarized Reasoning Chain
To achieve the Aligner module’s optimization goals, our motivation is to leverage a crucial reasoning chain related to the query-answer pair. This is especially important for multi-hop problems requiring intermediate reasoning steps to bridge information gaps. Incorporating such reasoning significantly improves answer quality, as demonstrated by methods like Chain-of-Thoughts (Wei et al., 2022) and O1 (Zhong et al., 2024). To explicitly identify the essential nodes within graph database, we propose an LLM-based summarization technique. By designing well-crafted prompts, we harness the capabilities of strong LLMs to generate comprehensive reasoning chains that logically connect the question to the answer. For example, Table 6 in the case study, given a query, “What is the name of the first Harry Potter novel?”, with the LLM-generated prior, we can identify the intermediate and critical node, J.K. Rowling. We believe this can serve as an excellent label. The deeper insights provided through reasoning are subsequently utilized in the dual node and representation alignment process to train the aligner module. More details for the summarization prompt are provided in experiments 5.10.
4.2.2. Node Alignment
The Aligner module utilizes the summarized reasoning chain to identify and align relevant nodes within the graph, effectively pruning redundant nodes and edges by filtering out unrelated information at the node level. Specifically, we employ a GNN (e.g., GraphTransformer (Shi et al., 2020) or GAT (Veličković et al., 2018)), to encode the structural information of the graph. The GNN produces node-level embeddings $\boldsymbol{n}_{g}$ based on the input graph:
(8) $\boldsymbol{n}_{g}=\text{GNN}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d},$
where $|\mathcal{V}|$ is the number of node and $d$ is feature dim. For the reasoning text, we employ SBERT to encode the textual description into a representation $\boldsymbol{r}_{s}$ that captures its semantic meaning:
(9) $\boldsymbol{r}_{s}=\text{SBERT}(t_{\text{reasoning}})\in\mathbb{R}^{d_{s}}$
We concatenate the embedding of each node in the subgraph with the query embedding. The concatenated embeddings are then passed through an MLP module, which generates a predicted importance score for each node. Finally, these scores are transformed into probability distributions using the softmax function, producing importance scores for both prediction and reasoning.
(10) $\boldsymbol{p}_{\text{prediction}}$ $=\text{Softmax}(\text{MLP}([\boldsymbol{n}_{g},\boldsymbol{q}^{\text{expand}}]))\in\mathbb{R}^{|\mathcal{V}|},$
(11) $\boldsymbol{p}_{\text{reasoning}}$ $=\text{Softmax}(\text{sim}(\boldsymbol{r}_{s},\boldsymbol{n}))\in\mathbb{R}^{|\mathcal{V}|}$
where $[,]$ means concat operation, $\boldsymbol{q}^{\text{expand}}$ is the query embedding broadcasted across all nodes. To align the node distribution, we minimize the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) between predicted probabilities $\boldsymbol{p}_{\text{prediction}}$ and the probabilities $\boldsymbol{p}_{\text{reasoning}}$. The KL divergence loss for a subgraph is given by:
(12) $\mathcal{L}_{\text{NA}}=\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\boldsymbol{p}_{\text{reasoning}}(i)\log\frac{\boldsymbol{p}_{\text{reasoning}}(i)}{\boldsymbol{p}_{\text{prediction}}(i)}$
Optimizing $\mathcal{L}_{\text{NA}}$ enables effective alignment of relevant knowledge.
4.2.3. Representation Alignment
To bridge the representation gap between graph structures and language-based descriptions, by treating the text representation derived from reasoning chain as the target label, Aligner module aligns the graph and text embeddings to encourage semantic consistency. We apply a mean pooling operation across the node embeddings $\boldsymbol{n}_{g}$ to obtain a unified graph-level representation $\boldsymbol{r}_{g}$:
(13) $\boldsymbol{r}_{g}=\text{POOL}(\boldsymbol{n}_{g})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\boldsymbol{n}_{g}(v)\in\mathbb{R}^{d}$
To unify the graph and text representations in a shared semantic space, we apply a contrastive loss with in-batch negative sampling. This loss encourages positive pairs (i.e., graph and text embeddings) to have higher similarity, while pushing apart non-matching pairs. Then, a shared-weight MLP layer is further designed to map $\boldsymbol{r}_{g}$ and $\boldsymbol{r}_{s}$ to dimension $d_{t}$ of LLM token embeddings:
(14) $\boldsymbol{\hat{r}}_{s}$ $=\text{MLP}(\boldsymbol{r}_{s})\in\mathbb{R}^{d_{t}},$
(15) $\boldsymbol{\hat{r}}_{g}$ $=\text{MLP}(\boldsymbol{r}_{g})\in\mathbb{R}^{d_{t}}$
The contrastive loss for representation alignment from $\boldsymbol{\hat{r}}_{g}$ to $\boldsymbol{\hat{r}}_{s}$ is defined as:
(16) $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s})=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{j})/\tau\right)}\right]$
where, $N$ is the batch size, $(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})$ is the $i$-th positive (graph-text) pair in the batch, $\tau$ is a temperature parameter to control the sharpness. Similarly, we can obtains the loss $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})$ from $\boldsymbol{\hat{r}}_{s}$ to $\boldsymbol{\hat{r}}_{g}$. The final representation alignment loss is obtained by:
(17) $\mathcal{L}_{RA}=\frac{1}{2}(\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})+\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s}))$
To achieve an optimized Graph Aligner, we perform joint optimization of node and representation Alignment. The total loss for the Graph Aligner is defined as:
(18) $\mathcal{L}_{\text{Aligner}}=\mathcal{L}_{\text{RA}}+\mathcal{L}_{\text{NA}}$
The parameters of GNN encoder are jointly optimized using the loss $\mathcal{L}_{\text{Aligner}}$ with a specified training step, which reflects degree of alignment. We evaluate the impact of the alignment degree in Figure 4 and the effectiveness in Section 5.7.
4.2.4. Graph Pruning and Representation Generation
We use the trained Graph Aligner to prune irrelevant nodes, and produce graph representations that are aligned with language. The Aligner performs graph pruning across the entire dataset. To start, we introduce a hyperparameter: the number of seed nodes, denoted as $n_{\text{seed}}$. For the graph’s node-level output, $\boldsymbol{n}_{\text{prediction}}$, the top $n_{\text{seed}}$ nodes are selected as seed nodes. Using the first-order neighbors of these seed nodes, we expand the adjacent nodes and edges. This process yields the newly constructed subgraph, denoted as $\mathcal{G}_{\text{Aligner}}=(\mathcal{V}_{\text{Aligner}},\mathcal{E}_{\text{Aligner}})$. What is more, we also utilize the trained Aligner across the entire dataset to produce the graph representation $\boldsymbol{r}_{g}$. The aligned subgraph $\mathcal{G}_{\text{Aligner}}$ serves as the input to LLM, enabling more efficient and accurate generation results. The efficiency of pruning is evaluated in section 5.8.
4.3. Graph Generator
The Aligner module enables efficient representation alignment and pruning of irrelevant knowledge, greatly facilitating integration with LLMs. To generate the final answer, we concatenate the text of graph $\mathcal{G}_{\text{Aligner}}$ with query tokens and input them into the LLM’s TextEmbedder ((The token embedding of LLM itself). This yields text tokens embedding: $\boldsymbol{r}_{t}=\text{TextEmbedder}([t_{q},\mathcal{G}_{\text{Aligner}}])\in\mathbb{R}^{(T_{q}+T_{g})\times d_{t}}$, where $T_{q}$ and $T_{g}$ mean the token length of query and graph. $\boldsymbol{r}_{t}$ is then concatenated with the previously learned graph embedding $\boldsymbol{r}_{g}$. The conditional probability of generating answer $Y$ given the aligned graph $\mathcal{G}_{\text{Aligner}}$ and the query $t_{q}$ is defined as:
(19) $p_{\theta}(Y|t_{q},\mathcal{G}_{\text{Aligner}})=\prod_{i=1}^{m}p_{\theta}(y_{i}|[\text{MLP}(\boldsymbol{r}_{g}),\boldsymbol{r}_{t}],y_{<i})$
To enhance efficiency of training and deployment, we fine-tune generator using parameter-efficient tuning method, like LoRA (Hu et al., 2021).
5. Experimental Settings
Table 1. Statistics of GraphQA Benchmark (He et al., 2024).
Dataset #Training #Validation #Test Task Type
ExplaGraphs 1,659 553 554 Commonsense reasoning
SceneGraphs 59,978 19,997 20,025 Visual scene understanding
WebQSP 2,826 245 1,628 Knowledge-based QA
5.1. Datasets and Metrics.
Following G-Retriever (He et al., 2024), we conducted experiments on the GraphQA benchmark (He et al., 2024), which includes ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). The dataset statistics are summarized in Table 1, with a data split of train:validation:test = 6:2:2.
  • ExplaGraphs is a dataset designed for commonsense reasoning and focuses on generating explanation graphs to predict stances in debates. It provides detailed, explicit, and commonsense-enriched graphs to evaluate arguments that either support or refute a given stance. The primary task is to determine whether an argument supports or opposes a stance.
  • SceneGraphs is a visual question answering dataset. Each graph provides detailed descriptions of objects, attributes, and relationships within an image. It is designed for tasks that involve spatial understanding and multi-step reasoning. The goal is to answer open-ended questions based on textual descriptions of scene graphs, with Accuracy used as the evaluation standard.
  • WebQSP is a large-scale, multi-hop knowledge graph question-answering dataset built on a subset of Freebase. It focuses on questions requiring multi-hop reasoning and includes facts within a two-hop neighborhood of the entities mentioned in the question. Since a single question may have multiple answers, the model’s performance is evaluated using F1, Hit@1, and Accuracy following Hu et al. (2024).
5.2. Baselines.
To assess the effectiveness of our proposed method, we compare it against four categories of baselines:
5.2.1. Inference-only.
This category includes approaches that employ frozen LLMs for question answering, using textual graph descriptions and queries as input. Different prompt strategies are considered, including:
  • Zero-shot: directly answering questions based on retrieved information,
  • Zero-CoT (Wei et al., 2022): enhances zero-shot reasoning by appending the phrase “Let’s think step by step.”, which encourages the model to explicitly generate intermediate reasoning steps.
  • CoT-BAG (Wang et al., 2024): adds “Let’s construct a graph with the nodes and edges first.” after providing the textual graph description,
  • KAPING (Baek et al., 2023): augments LLMs with knowledge graph facts at the prompt level, enabling more accurate QA without requiring model fine-tuning retrieving relevant graph triples.
  • ToG (Sun et al., [n. d.]): integrates LLMs with KGs to support deep and reliable reasoning through interactive exploration of entities and relations.
5.2.2. Raw Fine-tuning.
In this setting, following (He et al., 2024), we fine-tune the LLM via parameter-efficient methods, without applying advanced reranking techniques. Two widely used approaches are included:
  • Prompt Tuning (Lester et al., 2021): learns a set of continuous prompt embeddings that guide the frozen LLM during inference, enabling lightweight adaptation to specific tasks.
  • LoRA (Hu et al., 2021): introduces low-rank adaptation layers into transformer weights, achieving efficient fine-tuning with significantly reduced computational and memory costs.
5.2.3. Reranker-based.
These baselines employ reranking models to refine the ranking of candidate documents before passing them into the LLM. Representative methods include:
  • GTE (General Textual Embedding) (Li et al., 2023b): developed by Alibaba DAMO Academy, with two variants (gte-base, 109M parameters; gte-large, 335M parameters), trained on large-scale relevance pairs to improve retrieval accuracy,
  • BGE (BAAI General Embedding) (Xiao et al., 2024): cross-encoder models with two variants (bge-reranker-base, 278M parameters; bge-reranker-large, 560M parameters), optimized for retrieval-augmented generation, offering high accuracy at the cost of efficiency,
  • G-RAG and G-RAG-RL (Dong et al., 2024): reranker models that leverage GNNs to incorporate document connections and semantic cues from abstract meaning representation graphs for context-aware ranking.
5.2.4. GNN-based.
This category integrates GNN encoders with LLM embeddings to enhance graph reasoning, including:
  • GraphToken (Perozzi et al., 2024): encodes graph structures as explicit prompts for LLMs, improving performance on graph reasoning tasks,
  • GNP (Tian et al., 2024): a plug-and-play framework that applies a GNN encoder with cross-modality pooling to enrich LLMs with knowledge graph information,
  • G-Retriever (He et al., 2024): introduces the GraphQA benchmark and employs retrieval-augmented generation with soft prompting, improving graph-based question answering, comprehension, and reducing hallucinations,
  • GRAG (Hu et al., 2024): proposes a divide-and-conquer strategy for efficient textual subgraph retrieval and combines text and graph views within LLMs for graph context-aware generation.
5.3. Implementation Details.
We use GraphTransformer (Shi et al., 2020), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2016) as GNN encoders, and Llama-2-7b-hf, Llama-2-13b-hf and Llama-3.1-8b (Touvron et al., 2023) as generators. For the retrieval process, we use the same retrieval results obtained through the PCST algorithm across all baselines to ensure a fair comparison. Both reranker-based and GNN-based methods apply LoRA for fine-tuning, except for GraphToken, whose results are from G-Retriever (He et al., 2024). All methods are compared using the same training hyperparameters where applicable, ensuring fair comparability (e.g., GNN layers and LoRA rank). In the Aligner module, we explore two hyperparameters: alignment degree and the number of seed nodes ($n_{\text{seed}}$), with the analysis shown in Figure 7. To summarize the reasoning chain, we employ Llama-3.1-70B-Instruct (Dubey et al., 2024). The models were trained on the training set, with optimal hyperparameters and early stopping determined using the validation set. Results are reported on the test set. All experiments were run on two NVIDIA A100 GPUs (80GB each).
To implement baselines, for GNP (Tian et al., 2024), we implemented the Graph Neural Prompting module within our framework, including components such as the GNN encoder, cross-modality pooling, and domain projector. For G-RAG and G-RAG-RL (Dong et al., 2024), we adopted their ranking approach, combining cross-entropy loss with pairwise ranking loss. However, since treating documents as nodes was infeasible in our case, we instead treated entities as nodes and employed the same GNN encoder as our method. We implemented GTE (Li et al., 2023b) and BGE-reranker (Xiao et al., 2024) using their official open-source models. Nodes and triples are ranked by query similarity and, as in our method, fed into the LLM for generation. For the $\text{GRAG}_{PCST}$ model (Hu et al., 2024), we reproduced its experiments. However, to ensure a fair evaluation, we standardized the retrieval process by using the PCST method for retrieval. This allowed us to directly compare it with their graph encoder approach.
5.4. Main Results
Table 2. Performance comparison using the same retrieval settings for all methods. Llama-2-7b-hf is used as generator. The table reports the mean and standard deviation (std) results across three random seeds. For methods marked with $\dagger$’, we reproduce the results on WebQSP to report both F1 and Accuracy. Results for methods marked with $\ddagger$’ are taken directly from (He et al., 2024). The best results are highlighted in bold, and the second-best results are underlined. ‘Improvement $(\triangle)$’ represents the gain of Align-GRAG over the second-best baseline. ‘*’ indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. $\uparrow$: higher is better.
Method ExplaGraphs SceneGraphs WebQSP
[…]
Inference-only
[…]
Raw Fine-tuning
[…]
Reranker-based
[…]
GNN-based
[…]
Improvement $(\triangle)$ +2.12%* +1.39% +1.68% +1.88%* +3.58%*
In this section, we conduct extensive experiments on the ExplaGraphs, SceneGraphs, and WebQSP datasets, comparing the performance of our method with 16 baseline methods. All baselines are implemented using the same retrieved text, and the same backbones, Llama2-7b (Touvron et al., 2023) and GraphTransformer (Shi et al., 2020). As shown in Table 2, the overall results demonstrate the effectiveness of Align-GRAG, which achieves performance improvements across all metrics on three datasets. Notably, on WebQSP dataset, Align-GRAG achieves a remarkable improvement in Accuracy compared to the second-best method. This improvement can be attributed to the innovative design of Align-GRAG, particularly its alignment-based pruning and optimization strategies for graph structures, showcasing its potential in complex graph reasoning and knowledge-based question answering tasks.
Comparing with Inference-only methods. Inference-only methods rely solely on the reasoning capabilities of LLMs without incorporating task-specific optimization, which often leads to subpar performance. Specifically, the Zero-shot approach, which does not depend on any carefully constructed prompts, suffers from limited performance due to its inability to leverage task-specific knowledge. In contrast, ToG achieves superior performance by iteratively executing beam search on KG, enabling the discovery of promising reasoning paths. These results underscore ToG’s effectiveness within inference-only approaches.
Comparing with raw fine-tuning methods. LoRA significantly outperforms Prompt Tuning and also surpasses inference-only methods. This highlights the effectiveness of task-specific optimization in graph QA scenarios. However, it still lags behind reranker-based and GNN-based methods, indicating that more advanced techniques are required to fully exploit graph structure.
Comparing with reranker-based methods. gte-large achieves the second-best performance on both the ExplaGraphs and WebQSP datasets (F1 metric), outperforming other reranker approaches. However, it still falls significantly short of Align-GRAG. This demonstrates that simply performing node and triple reranking is insufficient to fully capture the graph structural information. Although G-RAG-RL also utilizes a GNN encoder to model graph structures, it relies on similarity with the query as a label, resulting in poor performance. This validates the effectiveness of our approach, which uses reasoning chains summarized by LLM as an optimization target. By aligning graph reasoning with LLM-derived reasoning chains, Align-GRAG achieves superior performance, showing that this alignment is crucial for graph-based reasoning tasks.
Comparing with GNN-based methods. G-Retriever ranks second on SceneGraphs, while GNP achieves the second-best Accuracy on WebQSP. However, both methods fall significantly short of Align-GRAG overall. This is because G-Retriever simply concatenates graph embeddings with LLM inputs using a projector, while GNP employs cross-modality pooling to fuse text and graph representations. Despite these efforts, their approaches remain limited. In contrast, Align-GRAG explicitly aligns the representation of graphs with language, effectively bridging the gap between these two spaces and achieving state-of-the-art results.
5.5. Ablation Study
Figure 3. Ablation study with different alignment strategy.
We perform an ablation study on WebQSP using three metrics to evaluate how the proposed modules contribute to performance improvement. As illustrated in Figure 3, we examine the following variants: (1) w/o Node Alignment: Removes the KL loss and knowledge pruning components. (2) w/o Representation Alignment: Excludes the contrastive loss and the graph embedding input to the LLM. (3) w/o Both: Removes both Node Alignment and Representation Alignment. (4) Random Alignment: Optimizes the aligner module using randomly generated labels. We observe that removing the Node Alignment module significantly lowers all evaluation metrics compared to the full model (Align-GRAG), indicating the effectiveness of node-level alignment optimization and node pruning. In contrast, while excluding the Representation Alignment module also leads to a performance decline, the impact is comparatively less significant. Performance further deteriorates when both modules are removed (w/o Both), highlighting the critical role of dual alignment between node and representation in filtering irrelevant knowledge and bridging the representation gap between the graph and language. The Random Alignment variant performs the worst, as random alignment fails to guide the model effectively. This emphasizes the necessity of meaningful alignment strategies and demonstrates that our LLM-summarized reasoning chain provides a highly useful label.
5.6. Generalization Analysis
Table 3. Generalization Analysis on different LLM and GNN backbones.
LLM Backbones Llama-2-7b-hf Llama-2-13b-hf Llama-3.1-8B
GNN Method ExplaGraphs WebQSP ExplaGraphs WebQSP ExplaGraphs WebQSP
[…]
GraphTransformer GNP 0.8704 0.7391 0.8880 0.7696 0.8892 0.7557
G-Retriever 0.8705 0.7366 0.9115 0.7739 0.9014 0.7482
Align-GRAG 0.8992 0.7626 0.9241 0.7789 0.9321 0.7712
GAT GNP 0.9061 0.7291 0.8989 0.7676 0.9062 0.7521
G-Retriever 0.7960 0.7414 0.8953 0.7737 0.9111 0.7719
Align-GRAG 0.9151 0.7309 0.9151 0.7573 0.9214 0.7742
GCN GNP 0.7545 0.7298 0.8682 0.7564 0.8572 0.7495
G-Retriever 0.8592 0.7352 0.9007 0.7521 0.8512 0.7527
Align-GRAG 0.8574 0.7377 0.9152 0.7592 0.8596 0.7573
In this section, we evaluate whether Align-GRAG generalizes across both GNN architectures and LLM scales on ExplaGraphs (Accuracy) and WebQSP (Hit@1) datasets. Table 3 summarizes results for three GNN backbones—GT (GraphTransformer) (Shi et al., 2020), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2016)—and three LLM families/sizes (Llama-2-7b-hf, Llama-2-13b-hf, Llama-3.1-8B), comparing against GNP and G-Retriever. We can observe that Align-GRAG attains the best score in 15/18 configurations, indicating robust generalization to diverse encoder–reasoner pairings.
What is more, we can find the strongest gains with GraphTransformer (GT). Using GT as the graph encoder yields the largest margins over GNP and G-Retriever on both datasets and across LLM sizes. This suggests GT’s capacity to model long-range dependencies synergizes with Align-GRAG’s dual alignment of nodes and representations, amplifying retrieval and reasoning quality.
Besides, benefits persist across attention-based GNNs. With GAT, Align-GRAG remains consistently competitive and often best, reinforcing that alignment mechanisms complement attention-driven neighborhood aggregation. Larger LLMs help, especially with GCN. Performance with 13B models generally surpasses 7B counterparts; the gap is most pronounced under the GCN backbone, where the LLM must compensate for weaker expressivity in the graph encoder. This indicates that larger LLMs better exploit aligned graph signals for multi-hop reasoning.
On the other hand, improvements hold on both ExplaGraphs and WebQSP datasets, demonstrating that Align-GRAG transfers well from explanation-centric graph tasks to question answering over knowledge graphs. Overall, Align-GRAG is architecture-agnostic yet particularly effective when paired with more expressive GNNs and larger LLMs, underscoring the importance of both high-capacity graph encoders and language models for learning fine-grained node-representation alignment.
5.7. Evaluation of Representation Alignment
(a) Align with Query
(b) Align with Summarization
(c) Align with Textualized Graph
Figure 4. Representation Alignment Analysis: The cosine similarity score between graph embeddings and language embeddings (aligned using the aligner module vs. the unaligned setting).
In this section, we evaluate whether the aligner can effectively bridge the representation gap? Specifically, we calculate the Cosine Similarity scores between the graph embeddings and the language embeddings of the query, summarization, and textualized graph, respectively, on test set. The graph embeddings include both the unaligned embeddings and the aligned embeddings (optimized using our contrastive loss). From the results in Figure 4, we observe that as the alignment degree increases, Cosine Similarity scores improve. For instance, similarity with summarization embeddings improves significantly as alignment progresses from 40 to 80 steps. For query and textualized graph embeddings, scores rise sharply between 0-4 steps. This indicates that training effectively reduces the representation gap between graph and language embeddings. After 8 steps, aligned embeddings consistently achieve higher similarity scores than unaligned across all text representations (query, summarization, and textualized graph), demonstrating the effectiveness of our approach. However, as shown in section 5.9.2, the alignment degree does not necessarily improve indefinitely. When it reaches a certain level, excessive alignment may over-alter the original graph information, compromising its accuracy.
5.8. Evaluating the Efficiency of Node Pruning
Table 4. Evaluation of Efficiency with and without node pruning. $\uparrow$ indicates results better than G-Retriever, while $\downarrow$ indicates worse results. (xx%) represents the percentage of tokens relative to G-Retriever. OOM means out of memory.
Method # Tokens Infer Time Hit@1
No Pruning Non-Retriever 100626.28 (387.02%) OOM -
BM25 (Robertson et al., 2009) 2569.57 (99.72%) 17:24 min 0.4394
Contriever (Izacard et al., [n. d.]) 2593.28 (100.01%) 17:37 min 0.4429
G-Retriever (Bienstock et al., 1993) 2576.66 (100.00%) 17:28 min 0.4502
Node Pruning with
Different Degree
Align-GRAG
[…]
In this section, we evaluate the efficiency improvements brought by node pruning in Aligner module. Specifically, we compare different retrieval methods in terms of average token consumption, average inference time, and Hit@1 scores on the test set of WebQSP. The w/ Aligner method is built on the initial retrieval results from PCST as the same as G-Retriever (He et al., 2024), where different seed nodes are selected to perform node pruning. The experiments were conducted in an inference-only setting. To fully utilize GPU resources and memory, we adjusted the batch size to be as large as possible under different average token conditions. For comparison, the BM25 and Contriever methods retrieve triples and ensure that the token consumption is roughly similar to that of G-Retriever. From Table 4, we observe that when the number of seed nodes ($n_{\text{seed}}$) is set to 6, our Aligner method achieves performance comparable to G-Retriever while only utilizing 27.11% tokens. This demonstrates the method’s ability to effectively prune irrelevant knowledge while preserving the essential information. However, when $n_{\text{seed}}$ is too small (e.g., $n_{\text{seed}}=4$), performance drops significantly, suggesting that some critical information may be pruned during the process. On the other hand, increasing the number of seed nodes to 15 (utilizing 60.04% of the tokens) leads to performance that significantly surpasses G-Retriever. This highlights the strength of our method in efficiently selecting useful knowledge and confirms that, in longer-context scenarios, removing irrelevant information enhances overall results. Additionally, we find that BM25 and Contriever both perform worse than G-Retriever when using a similar number of tokens. This suggests that the PCST retrieval method is better suited for graph databases, as it more effectively captures the connections between knowledge.
5.9. Impact of Hyperparameters
In this section, we analyze how hyperparameters affect performance, including: the Number of Seed Nodes $n_{\text{seed}}$, Align Degree (training step) and Top K retrieval.
5.9.1. Impact of Seed Nodes
We analyze the impact of the number of seed nodes ($n_{\text{seed}}$) on model performance. The experiments are conducted on the WebQSP dataset, where we evaluate the Hit@1, F1 and Accuracy metrics. From the experiments in Figure 5, we observe that the Hit@1, F1 and Accuracy performance peaks when the number of seed nodes is set to 25. Beyond this point (from 25 to 30), the performance starts to decline. This indicates that including too many nodes introduces a significant amount of irrelevant knowledge, which negatively impacts the model. Our pruning strategy effectively eliminates irrelevant knowledge to enhance model performance. On the other hand, when the number of seed nodes is as low as 5, the performance is considerably poor. This suggests that excessive pruning removes crucial knowledge, which is detrimental to performance. This highlights a trade-off: pruning reduces noise and improves performance, but over-pruning leads to the loss of essential knowledge.
Figure 5. Hyperparameters Analysis of the Number of seed nodes.
5.9.2. Impact of Align Degree
This section examines how the Align Degree (number of training steps for Aligner module) influences model performance. As shown in Figure 6, we evaluate the Hit@1, F1 and Accuracy metrics on the WebQSP dataset. From the experimental curve of Align Degree, we observe that the Hit@1, F1 and Accuracy metrics peak at around 60 epochs before declining. This indicates that, as training progresses, bridging the representation gap between the graph and language helps the LLM better understand graph data. However, excessive training may lead to overfitting, which disrupts graph information and ultimately causes a drop in performance. This suggests that there is an optimal point in training where the representation alignment is most effective, and training beyond this point can be detrimental.
Figure 6. Hyperparameters Analysis of Align degree.
5.9.3. Impact of Top K retrieval
Figure 7 illustrates the impact of varying the Top K retrieval of entities and relations on model performance across Hit@1, F1 and Accuracy. By analyzing the trends in the graphs, we can derive insights into the effect of Top K on model performance. All three metrics show a similar trend: performance improves as K increases, peaks at K = 10, and then declines. This suggests that K = 10 strikes the optimal balance between retrieving relevant results and avoiding noise. Smaller K values may miss relevant information, while larger K values dilute relevance, reducing precision and retrieval quality. These findings highlight the importance of selecting an appropriate K value to maximize performance in retrieval-based systems.
Figure 7. Effect of Top K retrieval.
5.10. Evaluating the Quality of Reasoning Chain
In this section, we conduct a fine-grained analysis of the quality of the reasoning chain by LLM summarization. This analysis is divided into two aspects. First, we evaluate Relevance and Faithfulness to assess whether the generated chains are aligned with the questions and answers. Second, we compare the QA performance of the naive-quality reasoning chain to highlight our method’s impact. What is more, a case study is used to gain a deeper understanding of the forms of summarization in Section 5.11.
5.10.1. Evaluation of Relevance and Faithfulness
To assess the quality of generated reasoning chain in terms of their relevance to question-answer context and their faithfulness to provided contextual information, we adopted LLM-as-a-judge framework (Gu et al., 2024). Specifically, we utilized GPT-4o (OpenAI, 2024), genimi-2.5-flash (Comanici et al., 2025) and human to perform the evaluation. A total of 100 samples were randomly selected for this analysis. The evaluation was conducted using a carefully designed prompt, as illustrated in Figure 10. The results of the experiment are presented in Figure 8. The analysis reveals that the summarized reasoning chains exhibit a high degree of relevance and faithfulness. Specifically, the summaries achieved relevance scores of 90%, 92%, and 92%, indicating that the majority of the generated content aligns well with the given question-answer context. Similarly, faithfulness scores of 87%, 89%, and 90% demonstrate that the summaries largely adhere to the factual information provided by the context, with minimal hallucination. Furthermore, the alignment between LLM evaluations and human evaluations indicates the reliability and rationality of LLM-based judge.
Figure 8. Evaluating Relevance and Faithfulness of Reasoning Chain.
5.10.2. Quantitative comparison
In this experiment, we evaluate the impact of reasoning quality on QA performance by comparing two approaches: our $\text{Reasoning}_{\text{Align-GRAG}}$, where both the question and answer are included during reasoning chain generation, and $\text{Reasoning}_{\text{Naive}}$, where only the question is used for reasoning chain generation. The results in Table 5 demonstrate that incorporating the answer during summary generation ($\text{Reasoning}_{\text{Align-GRAG}}$) consistently outperforms the naive approach ($\text{Reasoning}_{\text{Naive}}$) across all evaluation metrics of datasets. Specifically, for ExplaGraphs, the accuracy rises significantly from 0.8592 using a naive reasoning approach to 0.8992 when employing $\text{Reasoning}_{\text{Align-GRAG}}$. For SceneGraphs, the accuracy improves from 0.8404 to 0.8804, illustrating the advantage of our method in enhancing graph-based reasoning accuracy. For WebQSP, The F1 score of $\text{Reasoning}_{\text{Align-GRAG}}$ is 0.5445, which is higher than the 0.5209 achieved by $\text{Reasoning}_{\text{Naive}}$, indicating that our strategy improves the balance of precision and recall in QA tasks. Similarly, the Hit@1 value increases from 0.7444 to 0.7626, showing that our approach helps the model identify correct nodes more accurately in graph alignment. These findings confirm that our method can consistently generate high-quality reasoning chains, which in turn lead to significant performance improvements.
Table 5. Quantitative Comparison of Different Reasoning Quality.
Method ExplaGraphs SceneGraphs WebQSP
[…]
$\text{Reasoning}_{\text{Align-GRAG}}$ 0.8992 0.8804 0.5445 0.7626 0.5700
$\text{Reasoning}_{\text{Naive}}$ 0.8592 0.8404 0.5209 0.7444 0.5409
Table 6. The case for summarizing the reasoning chain. The text from SUMMARIZE_PROMPT, Question, Answer, and Graph DataBase are concatenated as input to generate the Response from LLM. We can see that LLM’s summarization accurately identified middle and critical node in graph, ‘J.K. Rowling’, for answer.
SUMMARIZE_PROMPT: You are a helpful assistant responsible for generating a comprehensive summary of the data provided below. Given question and answer, and related graph data base. Please concatenate all of these into a single, comprehensive description. The description should logically connect the question to the answer. Make sure to include information collected from all descriptions.
Question: what is the name of the first harry potter novel?
Answer: harry potter and the philosopher’s stone
Graph DataBase:
node_id,node_attr
0,harry potter and the chamber of secrets
1,harry potter and the philosopher’s stone
3,j. k. rowling
7,harry potter and the half-blood prince
9,harry potter and the prisoner of azkaban
11,harry potter and the goblet of fire
16,harry potter
24,harry potter and the deathly hallows
57,fiction
59,harry potter literary series
76,professor severus snape
98,fantasy
src,edge_attr,dst
16,freebase.equivalent_topic.equivalent_domain,91
91,freebase.domain_profile.featured_views,806
199,book.written_work.subjects,455
59,book.book_subject.works,199
3,book.author.works_written,670
24,media_common.adapted_work.adaptations,46
59,book.book_subject.works,371
0,book.book.characters,325
24,book.book.genre,57
24,book.book_edition.book,24
9,book.book.genre,98
190,book.book_edition.book,24
1,book.book.genre,224
24,book.book_edition.book,24
7,book.book.genre,98
59,book.literary_series.fictional_universe,16
478,book.book_edition.book,24
Summarized Reasoning Chain: J.K. Rowling wrote the first Harry Potter novel, “Harry Potter and the Philosopher’s Stone.”
5.11. Case study
In this case study, we investigate how summarization enhances the discovery of critical intermediate nodes within reasoning chains, as illustrated in Table reftab:summarization. Using well-crafted prompts in conjunction with a graph database, we show that concise, targeted summaries can surface pivotal entities—such as ‘J.K. Rowling’—that bridge the semantic gap between an initial query and the final answer. Our analysis evaluates the extent to which summarization reliably extracts these key connectors, thereby enabling coherent multi-hop reasoning that would otherwise be obscured within the raw graph structure. We further demonstrate that, without the interpretive layer provided by summarization, the underlying graph database lacks the necessary contextual signals to assemble the correct reasoning path end-to-end, underscoring the value of summarization as a catalyst for interpretable and effective knowledge retrieval.
5.12. LLM Prompts
We introduce three prompts: the Summarization Prompt, the Generator Prompt, and the Relevance and Faithfulness Prompt. Each serves a specific function in enhancing the performance and evaluation of LLMs. Below is a detailed explanation of their purposes and applications, as illustrated in Figure 9, Figure 11, and Figure 10.
Summarization Prompt You are a helpful assistant responsible for generating a comprehensive summary of the data provided below. Given question and answer, and related graph data base. Please concatenate all of these into a single, comprehensive description. The description should logically connect the question to the answer. Make sure to include information collected from all descriptions. Question: {question} Answer: {Answer} Graph DataBase: {Textualized Graph}
Figure 9. Prompt for Summarization.
The Summarization Prompt is specifically designed to generate a comprehensive and cohesive summary by synthesizing three key inputs: a given question, its corresponding answer, and the relevant textualized graph data. The goal of this prompt is to create a logical connection between the question and the answer, while also integrating information extracted from the provided graph data. This ensures the output not only answers the question but also demonstrates an understanding of the contextual relationships within the graph. By combining these elements, the prompt facilitates the generation of a rich, informative summary that serves as a bridge between structured and unstructured data.
Relevance and Faithfulness Prompt Evaluation of Relevance: Evaluate the relevance of the Reasoning Chain in answering the QUESTION. A relevant Reasoning Chain contains information that helps answer the question, even if partially. Return one of the following labels: ’Relevant’, or ’Irrelevant’ without any additional response. Evaluation of Faithfulness: Evaluate the following Reasoning Chain for faithfulness in answering the QUESTION. A faithful response should include information that helps answer the question, even if partially, avoid inventing new details, and not contradict the context. Return one of the following labels: ’Faithful’ or ’Not Faithful’ without any additional response.
Figure 10. Evaluation of Relevance and Faithfulness Prompt.
Generator Prompt Prompt for WebQSP and SceneGraphs datasets: Textualized Graph: {Textualized Graph}. Please answer the given question. Question: {question} Answer: Prompt for ExplaGraphs dataset: Textualized Graph: {Textualized Graph}. Argument 1: {arg1} Argument 2: {arg1} Question: Do argument 1 and argument 2 support or counter each other? Answer in one word in the form of ‘support’ or ‘counter’. Answer:
Figure 11. Prompt for Generator.
The Relevance and Faithfulness Prompt is an evaluation-focused tool that serves two distinct but complementary purposes:
  • Relevance Evaluation: This component assesses whether the reasoning chain provided is relevant to the given question. A relevant reasoning chain must contain information that contributes meaningfully to answering the question, even if the contribution is partial. The model is required to return one of two labels: Relevant or Irrelevant, with no additional commentary.
  • Faithfulness Evaluation: This component measures the faithfulness of the reasoning chain in responding to the question. A faithful reasoning chain should accurately represent the information in the context, avoid fabrications, and not contradict the provided data. Similarly, the model outputs one of two labels: Faithful or Not Faithful. This dual evaluation ensures the generated responses are both accurate and contextually grounded.
The Generator Prompt is tailored to the dataset being used, with its structure varying depending on the specific requirements of the dataset. For WebQSP and SceneGraphs, the prompt focuses on answering a question by leveraging the provided textualized graph data. This encourages the model to extract and synthesize relevant information from graph representations to deliver a concise and accurate response. For ExplaGraphs, the prompt shifts its focus to reasoning. It requires the model to evaluate the relationship between two arguments (e.g., whether they support or counter each other), based on the provided textualized graph data. The model outputs a single-word response, either support or counter, reflecting the nature of the relationship.
6. Conclusion
This work introduces Align-GRAG, a novel framework designed to address the core challenges in graph-based retrieval-augmented generation: the inclusion of irrelevant knowledge during retrieval and the representation gap between graph embeddings and language model representations. To tackle these issues, Align-GRAG incorporates an innovative graph aligner module that performs dual alignment. The Graph Node Alignment mechanism prunes irrelevant nodes, preserving critical structural information, while Graph Representation Alignment bridges the semantic gap, ensuring a unified representation space between graph and language models. Extensive experiments on the GraphQA benchmark demonstrate that Align-GRAG consistently outperforms existing baselines across multiple tasks, such as commonsense reasoning and knowledge graph understanding. By effectively integrating graph reasoning with LLM-based generation, Align-GRAG achieves more accurate, context-aware, and reliable outputs. This work paves the way for future advancements in GRAG systems, offering a robust solution for leveraging structured graph data in retrieval-augmented generation tasks.
7. Limitations
Despite the promising results demonstrated by Align-GRAG, this work has certain limitations. Due to resource constraints, we were unable to conduct experiments on larger LLMs, leaving the effectiveness of the proposed alignment approach on more powerful models uncertain. Additionally, since our method requires the generation and utilization of graph embeddings, it cannot be directly implemented on closed-source models, such as GPT-4, which restricts access to internal embedding representations. These limitations highlight potential areas for future exploration, such as validating the scalability of the approach with state-of-the-art LLMs and developing techniques to adapt Align-GRAG for closed-source environments.